Construction of initial data sets for Lorentzian manifolds with lightlike parallel spinors

Bernd Ammann^{∗}  Klaus Kröncke  Olaf Müller

September 20, 2021

Abstract

Lorentzian manifolds with parallel spinors are important objects of study in several branches of geometry, analysis and mathematical physics. Their Cauchy problem has recently been discussed by Baum, Leistner and Lischewski, who proved that the problem locally has a unique solution up to diffeomorphisms, provided that the intial data given on a space-like hypersurface satisfy some constraint equations. In this article we provide a method to solve these constraint equations. In particular, any curve (resp. closed curve) in the moduli space of Riemannian metrics on M with a parallel spinor gives rise to a solution of the constraint equations on M a, b (resp. M S^{1}).

Keywords: Parallel spinor, Lorentzian manifolds, Riemannian manifolds, special holonomy, moduli spaces

1 Introduction

1.1 Parallel spinors

Let \overline{N}, h̄ be a connected n 1 -dimensional oriented and time-oriented Lorentzian manifold with a fixed spin structure. The bundle of complex spinors will be denoted by Σ\overline{N}. The spinor bundle carries a natural Hermitian form , of split signature, a compatible connection, and Clifford multiplication. In the present article we search for manifolds with a parallel spinor, i.e. a (nontrivial) parallel section of Σ\overline{N}. Understanding Lorentzian manifolds with parallel spinors is interesting for several reasons.

The first reason is that Riemannian manifolds with parallel spinors provide interesting structures. We want to briefly sketch some of them: Parallel spinors provide a powerful technique to obtain Ricci-flat metrics on compact manifolds: All known closed Ricci-flat manifolds carry a non-vanishing parallel spinor on a finite covering. Parallel spinors are also linked to “stability”, defined in the sense that the given compact Ricci-flat metric cannot be deformed to a positive scalar curvature metric. This condition in turn is linked to dynamical stability of a Ricci-flat metric under Ricci flow: A compact Ricci-flat metric is dynamically stable under the Ricci flow if and only if it cannot be deformed to a positive scalar curvature metric ([20] and [23, Theorem 1.1]).

∗B.A. has been partially supported by SFB 1085 Higher Invariants, Regensburg, funded by the DFG.

Infinitesimal stability was proven for metrics with a parallel spinor in [39] and local stability (for manifolds with irreducible holonomy) in [15]. Manifolds with parallel spinors provide interesting moduli spaces, see [3]. Furthermore parallel spinors help to understand the space of metrics with non-negative scalar curvature. The stability property explained above implies that every homotopy class in the space of positive scalar curvature metrics which is known to be nontrivial also remains non-trivial in the space of metrics with non-negative scalar curvature, see [35]. It is interesting and challenging to see to which extent it is possible to find Lorentzian analogues to these results.

A second reason to be interested in parallel spinors on arbitrary semi-Rie-mannian manifolds is that their existence implies that the holonomy is special [27, 28], [11], [22], [13], [8]. Thus the construction of Lorentzian manifolds with parallel spinors provides examples of manifolds with special holonomy.

A third reason is that parallel spinors are relevant in many fields of theoretical physics. For example parallel spinors on Lorentzian manifolds are often viewed as generators for the odd symmetries in a supersymmetric theory (see e.g. [16]). Parallel spinors are also important in several physical theories of Kaluza-Klein type, i.e. involving additional compactified dimensions. Mathematically we model them by a semi-Riemannian submersion π T B from a high-dimensional Lorentzian manifold T to a macroscopically visible 3+1dimensional space-time B. In this context one has to make sure that spectral and other analytic properties of Dirac operators on T are comparable to the corresponding properties on B. This requires the existence of harmonic spinors on the fibers M π^{1} b , b B. If the scalar curvature assumed to be nonnegative, these harmonic spinors are parallel which in turn implies that the fibers are Ricci-flat. Varying b yields a family of metrics on M with parallel spinors. One thus obtains a map B od _{}M , where od _{}M is the (pre-)moduli space of metrics with a parallel spinor on some covering of M, which is also a central ingredient in this article.

1.2 The Cauchy problem for parallel light-like spinors

Important progress about Lorentzian manifolds with parallel spinors was recently achieved by H. Baum, T. Leistner and A. Lischewski [10, 30, 29], see also [9] for associated lecture notes. In particular, these authors showed the well-posedness of an associated Cauchy problem which we will now describe in more detail and which will be the main topic of the present article.

Let again \overline{N}, h̄ be a time- and space-oriented Lorentzian spin manifold, and let , be the Hermitian product on Σ\overline{N} with split signature. The Clifford multiplication on \overline{N}, h̄ will be denoted by .

Note that for any spinor φ on a Lorentzian manifold one defines the Lorentzian Dirac current V_{φ} Γ T \overline{N} of φ by requiring that

h̄ X, V_{φ} X φ, φ

holds for all X T \overline{N}. Recall that on Lorentzian manifolds the Clifford action of vector fields on spinors is symmetric, in contrast to the Riemannian case, where it is anti-symmetric.

Manifold Metric Dimension Type  Clifford  Scalar product
multiplication  on spinors
\overline{N}  h̄  n 1  Lorentzian  ,
N  h  n  Riemannian  ,
M  g  m n 1 Riemannian  ,
Submanifold Weingarten map
N \overline{N}  \overline{W}
M N  W

Figure 1: Notation for manifolds, Clifford multiplications and Weingarten maps

If φ is parallel, then V_{φ} is parallel as well. One can show that V_{φ} is a future oriented causal vector [9, Sec. 1.4.2, Prop. 2]. Thus V_{φ} is either time-like or light-like everywhere. In the time-like case, the Lorentzian manifold locally splits as a product N, h ℝ, dt^{2} , where N, h is a Riemannian manifold with a parallel spinor. So with respect to a suitable Cauchy hypersurface the understanding of such metrics directly relies on the corresponding results on Riemannian manifolds.

In this article we are concerned with the case, that φ is a parallel spinor with a light-like Dirac-current V_{φ}. This problem was studied in [10], [30], and [29].

Let N be a spacelike hypersurface in this Lorentzian manifold with induced metric h. For a future-oriented (time-like) normal vector field T with h̄ T, T 1 we define the Weingarten map \overline{W} ¯ T. We use the symbol \overline{W} instead of W, as the latter one will be used for the Weingarten map of hypersurfaces M in N, which will play a central role in our article. The Hermitian product , and the Clifford multiplication on Σ\overline{N} induce a positive definite Hermitian product and a Clifford multiplication on Σ\overline{N} _{N} which can be characterized by the formulas

φ, ψ T φ, ψ , X φ iT X φ

for all x N, φ, ψ Σ_{x}\overline{N}, X T_{x}N. The spin structure on \overline{N} induces a spin structure on the spacelike hypersurface N. Let ΣN be the associated spinor bundle of N. With standard tools about Clifford modules, one sees that there is a bundle monomorphism ι ΣN Σ\overline{N} _{N} over the identity of N such that the Clifford multiplication and the Hermitian product on ΣN are mapped to and , on Σ\overline{N} _{N}. The bundle monomorphism ι is a bundle isomorphism if and only if n is even. However note that ι is not parallel, i.e. the connection is not preserved under ι. More precisely ι is a linear pointwise expression in \overline{W}. From now on we identify ΣN with its image in Σ\overline{N} _{N} under ι, taking the connection from ΣN. As already mentioned above we will also consider hypersurfaces M of N and the spinor bundle ΣM of M. However the relation between spinors on M and on N is a bit easier than between spinors on \overline{N} and on N. One can work with an embedding ΣM into ΣN _{M} which preserves the scalar product , . We can even choose the embedding such that the Clifford multiplications coincide, however in the literature another embedding is often used which does no longer preserve the Clifford multiplication. In our application in Section 5 it depends on the parity of n which embedding is more convenient for us, see also Subsection 2.3 and Appendix B for further information about hypersurfaces and spinors. Thus we want to use for the Clifford multiplication on M in contrast to which is used for the Clifford multiplication on N.

For any spinor φ Γ ΣN we associate — see e.g. [29, (1.7)] — its Riemannian Dirac current U_{φ} Γ T N by requiring

h U_{φ}, X i X φ, φ ,  X T N.  (1)

By Lemma 20 in Appendix A we see, that the spinor φ satisfies h U_{φ}, U_{φ} φ ^{4} if and only if we have for any p N

iφ p V φ V T_{p}N .  (2)

Now, if we assume that \overline{N}, h̄ carries a non-vanishing parallel spinor, then this spinor induces a spinor φ on N, h such that

_{X}^{N}φ \frac{i}{2}\overline{W} X φ,  X T N,  (3)
U_{φ} φ iu_{φ}φ,  (4)

where U_{φ} Γ T N is defined as above and where

u_{φ}  \overline{h U_{φ}, U_{φ}},  (5)

see [10, (4) and following] or [29, (1.6) and (1.8)].

Note that the constraint equations (1)–(5) are not independent from each others. We comment on this in Appendix A.

Conversely, if N, h is given as an abstract Riemannian spin manifold, and if \overline{W} Γ End T N , and φ Γ ΣN satisfy (2), (3) and (4) with U_{φ} and u_{φ} defined by (1) and (5), then there is a Lorentzian manifold \overline{N}, h̄ with N as a Cauchy hypersurface and a parallel spinor, such that h is the induced Riemannian metric, \overline{W} the Weingarten map and φ is induced by a parallel spinor on \overline{N}, h̄ , see [29, Consequence of Theorems 2 and 3]. This was proven by solving the associated wave equations by using the technique of symmetric hyperbolic systems. A simpler approach, going back to a remark by P. Chrusciel was later given in [36, Chap. 4].

The question arises on how to solve these constraint equations. In the present article we will describe a new method to obtain solutions of these constraint equations. We will see how any smooth curve in the (pre-)moduli space of closed m n 1 -dimensional Riemannian manifolds with a parallel spinor together with a scaling function yields a solution to the constraint equations, see our Main Construction 15 in Section 6. And thus the well-definedness of the Cauchy problem implies the existence of an associated n 1 -dimensional Lorentzian with a parallel spinor. Such a relation between families of metrics with special holonomy and solutions of the constraint equations was already conjectured by Leistner and Lischewski, see [29]. We essentially show that the conditions in [29, Table 1] is satisfied if and only if the divergence condition (31) in our Appendix D is satisfied.

We also derive versions of this construction to obtain initial data on a compact Cauchy hypersurface (without boundary). A first idea is to use a smooth closed curve in the (pre-)moduli space of closed m-dimensional Riemannian manifolds with a parallel spinor together with a scaling function. However this will in general not lead to a solution of the original constraint equation, but to a twisted version thereof, see Main Construction 17. This will lead to a Lorentzian manifold with a parallel twisted spinor. Here the twist bundle is always a complex line bundle with a flat connection.

In special cases – more precisely assuming the “fitting condition” introduced in Section 6 we however obtain solutions of the constraint equations in the original (i.e. untwisted) sense, see Main Construction 16.

We consider it as remarkable that compared to other classical diffeomorphism invariant Cauchy problems in Lorentzian geometry, e.g. the Cauchy problem for Ricci-flat metrics, we get a large quantity of solutions to the constraint equations. Furthermore it is amazing that the solutions in our situation correspond to curves in the moduli space, while the set of solutions of the constraints in classical problems have no similar description. An important input for our article was the smooth manifold structure of the premoduli space od _{}M and the fact that the BBGM connection preserves parallel spinors along divergence free Ricci-flat families of metrics. Ricci-flat deformations of a metric thus preserve the dimension of the space of parallel spinors. The infinitesimal version of this should be seen as some kind of Hodge theory: infinitesimal Ricci-flat deformations can be viewed as elements of ker D^{T^{}M} ^{2} while infinitesimal deformations with parallel spinors can be viewed as elements of ker D^{T^{}M} — and on compact manifolds Hodge theory tells us that ker D^{T^{}M} ^{2} ker D^{T^{}M}. Let us also compare the results of this paper to the tightly related recent preprint [2]. While the current paper constructs initial data for Lorentzian manifolds with a parallel spinor, one of the major themes in [2] is to get topological obstructions for a closed spin manifold to be a spacelike hypersurface in a Lorentzian manifold with a parallel spinor. We also mentioned that the special case of 3 1 -dimensional Lorentzian manifolds (i.e. initial data on 3-dimensional manifolds) was studied in the recent preprint [32].

The structure of the article is as follows. In Section 2 we fix some conventions, recall and extend known facts about Clifford modules and about spinors on hypersurfaces. We also explain how to differentiate a spinorial expression such as the Levi-Civita derivative ^{g}φ of a spinor φ with respect to the Riemannian metric g. The known results for defining this differential along a path of metrics is recalled in Subsection 2.4, this includes the BBGM parallel transport and the associated connection which arises from the universal spinor bundle construction; the concrete calculations are a central tool of the article and are carried out in Section 3. In Section 4 we show that for divergence free Ricci-flat deformations, the BBGMconnection preserves parallel spinors, a result which we consider an interesting outcome of the article independently of the application to the constraint equations, mainly discussed in the article. In Section 5 we use this construction to obtain solutions of the constraint equations (1) – (5). In Section 6 we establish the connection to curves in (pre-)moduli spaces and we also discuss on how to obtain solutions on compact Cauchy hypersurfaces. The article ends with several appendices where we provide some details about facts which are already well-known, but where adequate literature was not available. We hope that these appendices help to make the article sufficiently self-contained.

Acknowledgements. We thank Helga Baum, Thomas Leistner and Andree Lischewski for bringing our attention to this problem and for enlightening talks and discussions. Our special thank goes to Andree Lischewski for sharing the above mentioned conjecture with us, already at an early stage. We also thank both referees for many good and substantial comments.

2 Preliminaries

2.1 Conventions

All Hermitian scalar products in this article are complex linear in the first entry and complex anti-linear in the second one. Let E be a Riemannian vector bundle over a compact Riemannian manifold M, g equipped with a metric connection and k ℕ_{0}. Then the Sobolev norm H^{k} of a section s Γ E is

s _{H^{k}} _{l 0}^{k} _{M} ^{l}s ^{2}dV ^{1 2} .

where dV is the volume element of g. As usual, we denote L^{2} H^{0}. We will write H^{k} g if we wish to emphasise the dependence of the norm on g.

We use the symbol ^{2} for the symmetrized tensor product, i.e. for a finitedimensional real vector space V the space of symmetric bilinear forms V V ℝ is denoted by ^{2} V^{}. Let SV^{}^{2} V^{}be the subset of positive definite symmetric forms on V. Applying this fiberwise to the tangent bundle T M we obtain the vector bundle ^{2} T^{}M and the bundle ST^{}M SM.

Definition 1. On a Riemannian manifold M the Einstein operator is the elliptic differential operator Δ_{E} Γ ^{2} T^{}M Γ ^{2} T^{}M , given by

Δ_{E}h ^{}h 2R̊h,  R̊h X, Y  h R_{e_{i},X}Y, e_{i}

where e_{1}, … e_{n} is a local orthonormal frame. Here, the curvature tensor is defined with the sign convention such that R_{X,Y} Z _{X,Y}^{2} Z _{Y,X}^{2}Z.

The Einstein operator is linked to the deformation theory of Ricci-flat metrics as follows: Let g be a Ricci-flat metric and g_{s} a smooth family of Ricci-flat metrics through g_{0} g, then

\frac{d}{ds} _{s 0}g_{s} h _{X}g

where X is a vector field and h ker div_{g} ker Δ_{E} , i.e. the essential part of a Ricci-flat deformation is an element in ker Δ_{E} . In particular, if g_{s} is a family of metrics with a parallel spinor, the essential part of its s-derivative is contained in ker Δ_{E} . For more details on the deformation theory of Ricci-flat metrics, see [12, Chapter 12 D].

2.2 Some facts about Clifford modules

In this subsection we briefly summarize some facts about representations of Clifford algebras. The interested reader might consult the first Chapter of the book by Lawson and Michelsohn [26] for further details.

Let E_{1}, …, E_{n} be the canonical basis of ℝ^{n}. The complexified Clifford algebra for ℝ^{n} with the Euclidean scalar product will be denoted by ℂl_{n}. In this section all Clifford multiplications are written as , independently on n; we use the convention that does not depend on n. In the following section we will then explain that the associated bundle construction then turns this Clifford multiplication both into the Clifford multiplication on M and the Clifford multiplication on N.

We define the complex volume element ω_{n}^{ℂ} as

ω_{n}^{ℂ} ϵ_{n}E_{1} E_{2} … E_{n} ℂl_{n},

where ϵ_{4k} ϵ_{4k 3} 1, ϵ_{4k 1} i and ϵ_{4k 2} i. The choices imply ω_{n}^{ℂ} ^{2} 1. The choice of sign for ϵ_{n} in the literature varies between different sources, our particular and unconventional choice yields an easy formulation of Lemma 2.

For n even, there is a unique irreducible complex represention of ℂl_{n}. Here and in the following “unique” will always “unique up to isomorphism of representation”. This unique representation will be denoted by Σ_{n}. It comes with a grading Σ_{n} Σ_{n}^{}Σ_{n}^{}given by the eigenvalues 1 of ω_{n}^{ℂ}.

For n odd there are two irreducible representations, and as ω_{n}^{ℂ} is central, Schur’s Lemma implies that ω_{n}^{ℂ} acts either as the identity or minus the identity which allows us to distinguish the two representations. For n odd, we will assume that Σ_{n} is the irreducible representation on which ω_{n}^{ℂ} acts as the identity. The other representation is called Σ_{n}^{#}. In the following Σ_{n}^{#} denotes either Σ_{n} or Σ_{n}^{#} for odd n and Σ_{n} for even n.

The standard inclusion ℝ^{n 1} ℝ^{n}, x x, 0 induces an inclusion ℂl_{n 1} ℂl_{n}. This turns Σ_{n} into a ℂl_{n 1}-module. In particular, the Clifford action of X ℝ^{n 1} is the same if X is viewed as an element of ℂl_{n 1} or ℂl_{n}. ^{1}

Thus, if n is odd, we can choose (and fix from now on) isometric isomorphisms of ℂl_{n 1}-modules J_{n} Σ_{n 1} Σ_{n} and J_{n}^{#} Σ_{n 1} Σ_{n}^{#}. If n is even, then we can choose (and fix from now on) an isometric isomorphism of ℂl_{n 1}-modules I_{n} Σ_{n 1} Σ_{n 1}^{#} Σ_{n}.

Lemma 2. For n odd, J_{n} Σ_{n 1} Σ_{n} and J_{n}^{#} Σ_{n 1} Σ_{n}^{#} are ℂl_{n 1}-linear isomorphisms of complex Spin n 1 -representations and they satisfy

e_{n} J_{n} φ  iJ_{n} ω_{n 1}^{ℂ}φ ,  (6)
e_{n} J_{n}^{#} φ  iJ_{n}^{#} ω_{n 1}^{ℂ}φ .  (7)

for all φ Σ_{n 1}. For n even there are isometric monomorphisms of complex Spin n 1 -representations J_{n}^{}Σ_{n 1} Σ_{n} and J_{n}^{#,} Σ_{n 1}^{#} Σ_{n}, satisying

e_{n} J_{n}^{}φ  iJ_{n}^{}φ ,  (8)
e_{n} J_{n}^{#,} φ  iJ_{n}^{#,} φ .  (9)

for all φ Σ_{n 1}^{#} and for any , .

Proof. If n is odd, then we see that ϵ_{n} iϵ_{n 1}, thus ω_{n}^{ℂ} iω_{n 1}^{ℂ} e_{n} ie_{n} ω_{n 1}^{ℂ}. This implies e_{n} ω_{n}^{ℂ} iω_{n 1}^{ℂ}. Hence, we obtain (6):

e_{n} J_{n} φ e_{n} ω_{n}^{ℂ} J_{n} φ iω_{n 1}^{ℂ}J_{n} φ J_{n} iω_{n 1}^{ℂ}φ .

1Here we do not follow the convention chosen by Bär, Gauduchon and Moroianu [6], where vectors in ℝ^{n 1} have two different Clifford actions on Σ_{n}, depending on whether it is seen as an element in ℂl_{n 1} or ℂl_{n}-action.

The verification of (7) is analogous. As these morphisms are ℂl_{n 1}-linear, and as Spin n 1 is contained in ℂl_{n 1}, the morphisms are also Spin n 1 -equivariant.

If n is even, then we define

J_{n}^{# ,} φ \frac{1}{\overline{2}} I_{n} φ ie_{n} I_{n} φ .

Equations (8) and (9) are obvious. The morphisms are no longer ℂl_{n 1}-linear. However e_{n} commutes with ℂl_{n 1} _{0} which is defined as the even part of ℂl_{n 1}, i.e. the sub-algebra generated by elements of the form X Y with X, Y ℝ^{n 1}. As a consequence the morphisms are ℂl_{n 1} _{0}-linear. As Spin n 1 is contained in ℂl_{n 1} _{0}, the morphisms are Spin n 1 -equivariant.

To show that the morphisms J_{n}^{# ,} are isometric (and thus also injective) it suffices to show I_{n} φ ie_{n} I_{n} φ , where φ Σ_{n 1} or φ Σ_{n 1}^{#}. We argue only for φ Σ_{n 1}, as the other case is completely analogous. As ω_{n 1}^{ℂ} and e_{n} anticommute, ie_{n} I_{n} φ is an eigenvector of ω_{n 1}^{ℂ} to the eigenvalue 1, while I_{n} φ is an eigenvector to the eigenvalue 1. As ω_{n 1}^{ℂ} is self-adjoint, orthogonality follows.

Remark 3. Let n be even. Clifford multiplication with e_{n} anticommutes with all odd elements in ℂl_{n 1}, in particular with ω_{n 1}^{ℂ} and vectors ℝ^{n 1} ℂl_{n 1}. Thus, the map φ I_{n}^{1} e_{n} I_{n} φ restricts to a vector space isomorphism A Σ_{n 1} Σ_{n 1}^{#} which anticommutes with Clifford multiplication by vectors in ℝ^{n 1}.

2.3 Spinors on hypersurfaces

In this section we want to describe how one can restrict a spinor on an ndimensional Riemannian spin manifold N, h to a spinor on an oriented hypersurface M carrying the induced metric g. As, this restriction is local, we can assume — by restricting to a tubular neighborhood of M and using Fermi coordinates, i.e. normal coordinates in normal directions — that N M a, b and h g_{s} ds^{2} where s a, b .

Conversely, given a family of metrics g_{s} and spinors φ_{s} Γ Σ^{g_{s}}M , s a, b , we want to obtain a spinor on N M a, b , h g_{s} ds^{2}.

Note that this description does not include the passage from a Lorentzian manifold to a spacelike hypersurface and vice versa, described in the introduction, which plays an important role in the work of Baum, Leistner and Lischewski. For this Lorentzian version analogous techniques are well presented in the lecture notes [9, Sec. 1.5].

For the Riemannian case, which is the goal of this section, a standard reference is [6]. However for the purpose of our article it seems to be more efficient to choose some other convention, following e.g. [1, Sec. 5.3]. To have a fluently readable summary, we do not include detailed proofs in this section. For self-containedness we include them in Appendix B.

In the following let T_{M}N be the vertical bundle of the projection N a, b , i.e. T_{M}N _{s a,b} T M s , viewed as a bundle over N.

For each s I a, b let P_{SO} M, g_{s} be the SO n 1 -principal bundle of positively oriented orthornormal frames on M, g_{s} , and let P_{SO} N be defined as the corresponding SO n -bundle over N, h . We note ν ∂ ∂s. The union of the bundles P_{SO} M, g_{s} is an SO n 1 -principal bundle over N M I, which we will denoted by P_{SO,M} N N (whose topology and bundle structure is induced by the SO n 1 -reduction of the GL n 1 - principal bundle of all frames over N containing ν). We get a map P_{SO,M} N P_{SO} N , mapping e_{1}, …, e_{n 1} to e_{1}, …, e_{n 1}, ν which yields an isomorphism P_{SO} N P_{SO,M} N _{SO n 1} SO n . Any topological spin structure on M yields a Spin n 1 -principal bundle P_{Spin,M} N with a Spin n 1 SO n 1 - equivariant map P_{Spin,M} N P_{SO,M} N in the usual way. This map induces

P_{Spin} N P_{Spin,M} N _{Spin n 1} Spin n P_{SO} N P_{SO,M} N _{SO n 1} SO n

which is a spin structure on N M I. In particular this induces a bijection from the set of spin structures on M (up to isomorphism) to spin structures on N (up to isomorphism).

The complex volume elements ω_{n 1}^{ℂ} ℂl_{n 1} and ω_{n}^{ℂ} ℂl_{n} then provide associated complex volume elements

ω_{M}^{ℂ} ϵ_{n 1}e_{1} e_{2} … e_{n 1} ℂl T M and ω_{N}^{ℂ} ϵ_{n}e_{1} e_{2} … e_{n} ℂl T N ,

where e_{1}, e_{2}, … is any positively oriented basis in T M resp. T N.

Using the associated bundle construction, we obtain for n odd:

Σ_{M}N P_{Spin,M} N _{Spin n 1} Σ_{n 1} _{s a,b} Σ^{g_{s}}M,

ΣN P_{Spin} N _{Spin n} Σ_{n},

Σ^{#}N P_{Spin} N _{Spin n} Σ_{n}^{#}.

The grading Σ_{n 1} Σ_{n 1}^{}Σ_{n 1}^{}yields a grading Σ_{M}N Σ_{M}N ^{}Σ_{M}N ^{} which is the grading by eigenvalues 1 of ω_{M}^{ℂ}.

For n even we have

Σ_{M}N P_{Spin,M} N _{Spin n 1} Σ_{n 1} _{s a,b} Σ^{g_{s}}M,

Σ_{M}^{#}N P_{Spin,M} N _{Spin n 1} Σ_{n 1}^{#} _{s a,b} Σ^{#,g_{s}}M,

ΣN P_{Spin} N _{Spin n} Σ_{n},

and the grading Σ_{n} Σ_{n}^{}Σ_{n}^{}yields a grading ΣN ΣN ^{}ΣN ^{}, given by eigenvalues 1 of ω_{N}^{ℂ}.

The Clifford multiplications ℝ^{n 1} Σ_{n 1}^{#} Σ_{n 1}^{#}, ℝ^{n} Σ_{n}^{#} Σ_{n}^{#}, induce Clifford multiplications T_{M}N Σ_{M}^{#}N Σ_{M}^{#}N, T N Σ^{#}N Σ^{#}N. As indicated before we use the symbols “ ” and “ ” to distinguish properly between the Clifford multiplication of M and the one of N.

Furthermore the maps J_{n}, J_{n}^{#}, J_{n}^{}, and J_{n}^{#,} from Lemma 2 induce vector bundle maps over id N N which are isometric injective ℂ-linear maps in each fiber. For n odd, we obtain fiberwise isomorphisms

J Σ_{M}N ΣN and J^{#} Σ_{M}N Σ^{#}N.

These maps commute with Clifford multiplication by vectors tangent to M. They satisfy

ν J φ  iJ_{n} ω_{M}^{ℂ} φ ,  (10)
ν J^{#} φ  iJ^{#} ω_{M}^{ℂ} φ .  (11)

for all φ Σ_{M}N.

On the other hand we get for n even

J^{}Σ_{M}N ΣN and J^{#,} Σ_{M}^{#}N ΣN.

These maps do not commute with Clifford multiplication by vectors tangent to M, and they are not surjective in any fiber. However, they satisfy

ν J^{}φ  iJ^{}φ ,  (12)
ν J^{#,} φ  iJ^{#,} φ .  (13)

for all φ Σ_{M}N and for any , .

The bundle maps defined above do not preserve the (partially defined) Levi-Civita connections on the bundles. They modify the connection by terms depending on the second fundamental form of M s in N. This is made precise in the following lemma.

Lemma 4. Let J^{# ,} be one of the bundle maps defined above. For X Γ T_{M}N , and φ Γ Σ_{M}N resp. Γ Σ_{M}^{#}N we have

_{X}^{N}J^{# ,} φ J^{# ,} _{X}^{M,g_{s}}φ \frac{1}{2}W X ν J^{# ,} φ ,  (14)

where the Weingarten map W is defined through g W X , Y ν _{X}^{N}Y _{X}^{M}Y.

A proof will be given in Appendix B.

In combination with equations (10), (11), (12), and (13) we will see later in this article that families of M, g_{s} -parallel spinors will lead to solutions of the generalized imaginary Killing spinor equation (3) which is one of the constraint equations.

2.4 BBGM connection

Let M be a compact spin manifold. We denote the space of all Riemannian metrics on M by . For every metric g we define

g Γ Σ^{g}M ,

and the disjoint union

_{g} _{g}.

One can equip and naturally with the structure of a Fréchet bundle π^{}. This bundle structure is needed to define a connection on this bundle. The oldest references that we used are by Bourguignon and Gauduchon [14] resp. by Bär, Gauduchon and Moroianu [6], this is why we choose the abbreviation BBGM for the four names. However we were told that there was also work by Bismut. The concepts were later properly formalized under the name ’universal spinor bundle’ in [4] and [31], where a finite-dimensional fiber bundle with a partial connection is constructed whose sections correspond to the elements of such that the parallel transport corresponds to the BBGM parallel transport. The connection is given in terms of horizontal spaces _{g,φ}, i.e. vector spaces satisfying

T_{g,φ} _{g,φ} Γ Σ^{g}M

in the sense of topological vector spaces such that

dπ^{}_{g,φ} T_{g,φ} T_{g} Γ ^{2} T^{}M

restricts to an isomorphism I_{g,φ} _{g,φ} Γ ^{2} T^{}M . In other words we obtain an injective map of Fréchet spaces L_{g,φ} Γ ^{2} T^{}M T_{g,φ} by postcomposing I_{g,φ} ^{1} with the inclusion of _{g,φ} T_{g,φ} .

The space _{g,φ} will smoothly depend on g, φ and will be compatible with the vector bundle structure of . We do not require more knowledge about Fréchet manifolds in our article, so we do not introduce this Fréchet structure in more detail.

To describe the horizontal space precisely we give the maps L_{g,φ}: We assume that h T_{g} Γ ^{2} T^{}M . Let g_{t} be a smooth path of metrics such that g_{0} g and \frac{d}{dt} _{t 0}g_{t} h. We consider the cylinder Z M 0, 1 with the metric g_{t} dt^{2}. Then, similar to [6] we can extend the spinor φ_{0} φ, defined on M 0 , uniquely to a spinor φ_{}on M 0, 1 satisfying _{∂_{t}}φ_{}0 where denotes the Levi-Civita connection on the cylinder. Note that φ_{}may be interpreted as a family, parametrized over 0, 1 , of spinor fields φ_{t} with respect to the family of metrics g_{t}. The restriction of φ_{}to M t , denoted by φ_{t} is a spinor for the metric g_{t}, and t φ_{t} is a smooth path in . We now define

L_{g,φ} h \frac{d}{dt} _{t 0}φ_{t}.

Lemma 5. The map L_{g,φ} is well-defined, linear, and smooth in g, φ .

The lemma follows from the construction of the universal spinor bundle given in [31]. The strategy in that paper is as follows: Let π^{S} SM M be the bundle of positive definite symmetric bilinear forms on T M. A complex vector bundle π^{univ} ΣM SM is constructed, called the universal spinor bundle, which carries a scalar product, a Clifford multiplication with vectors in T M and partial connection on ΣM with respect to π^{S} SM M. The Clifford multiplication is given by a bilinear map cl_{g} T_{π^{S} g}M Σ_{g}M Σ_{g}M for every g SM. Note that g is a scalar product on one single tangent space, namely on T_{π^{S} g}M, and thus Σ_{g} is thus a fiber of the vector bundle π^{univ} ΣM SM. This Clifford multiplication shall satisfy the Clifford relations and shall depend smoothly on g. By “partial connection” we mean that for any section φ of the bundle π^{univ} ΣM SM, the covariant derivative _{X}φ is defined for some X T SM if and only dπ^{S} X 0 (i.e., X is vertical for π^{S}). This partial connection comes from the vertical connection defined in [31, Definition 2.10], and allows to define a map L_{g,φ} as above. In particular this shows that \frac{d}{dt} _{t 0}φ_{t} does not depend on how we choose g_{t}, but only its derivative a t 0.

3 Variation of the parallel spinor equation

Let us first recall a result from [4]: For g, φ  let κ̂_{g,φ}  _{g,φ} Γ Σ^{g}M

T^{}M be the map defined by

κ̂_{g,φ} h  \frac{1}{4} _{i j} _{e_{i}}^{g}h , e_{j} e_{i} e_{j} φ.  (15)

In [4, Sec. 4.3] a similar map was defined, namely κ_{g,φ} T_{g,φ} _{g,φ} Γ Σ^{g}M Γ Σ^{g}M T^{}M which is related to κ̂_{g,φ} by the formula κ_{g,φ} h, ψ κ̂_{g,φ} h _{}^{g}ψ. It was shown in [4, Lemma 4.12] that this is related to the differential of the map K_{X} , g, φ g, _{X}^{g}φ as follows:

d_{g,φ}K_{X} h, ψ h, κ_{g,φ} h, ψ X  h, κ̂_{g,φ} h X _{X}^{g}ψ .  (16)

We define the Wang map _{g,φ} as the composition

2 T^{}M ^{2} T^{}M ^{cl^{g} ,φ id} Σ^{g}M T^{}M

so that

g,φ h X _{j 1}^{n} h X, e_{j} e_{j} φ .

It is easy to see (c. f. [15, Lemma 2.3, 1.]) that

g,φ h φ h

and

X _{g,φ} h _{g,φ} _{X}h _{g, _{X}φ} h .

Lemma 6. If φ is a parallel spinor on M, g , then the diagram

2 T^{}M Γ Σ^{g}M T^{}M Γ Σ^{g}M T^{}M_{g,φ} 4κ̂_{g,φ} φ div^{g} D^{T^{}M}

commutes. In other words we have for any h ^{2} T^{}M

D^{T^{}M} _{g,φ} h 4κ̂_{g,φ} h φ div^{g} h ,

where, for a vector bundle E over M with connection we define the Dirac operator D^{E} on Σ^{g}M E by cl^{g} ^{Σ^{g}M E} where ^{Σ^{g}M E} is the product connection.

Proof. Let h Γ ^{2} T^{}M . Then _{g,φ} h _{j 1}^{n} h , e_{j} e_{j} φ. We now assume that e_{1}, …, e_{n} is a local orthonormal frame satisfying e_{j} _{p} 0 at p M for 1 j n, and we calculate in p

D^{T^{}M} _{g,φ} h  _{i 1}^{n} e_{i} _{e_{i}}  _{g,φ} h

_{i,j 1}^{n} _{e_{i}}h , e_{j} e_{i} e_{j} φ

4κ̂_{g,φ} h _{i 1}^{n} _{e_{i}}h , e_{i} e_{i} e_{i} φ

4κ̂_{g,φ} h φ \begin{matrix}_{i 1}^{n} _{e_{i}}h , e_{i} \\ \end{matrix}_{div^{g} h} .

Very similarly we prove for arbitrary sections φ and h:

\begin{aligned}D^{T^{}M} _{g,φ} h  X & 4κ̂_{g,φ} h X div^{g} h X φ \\ & 2 _{h^{#} X}φ h^{#} X D^{g}φ, \end{aligned}  (17)

where h^{#} End T M is defined as h^{#} X h , X ^{#}.

The following lemma is a straightforward generalization of a Lemma by McKenzie Wang [39, Lemma 3.3 (d) for k 1], see also [15, Prop. 2.4].

Lemma 7. For φ Γ Σ^{g}M and h Γ ^{2} T^{}M we have

D^{T^{}M} ^{2} _{g,φ} h  _{g,φ} Δ_{E}h 2 _{ℓ} _{g, _{e_{ℓ}} φ} _{e_{ℓ}}h _{g,D^{2}φ} h

2 X _{ℓ} h^{}e_{ℓ} R_{X,e_{ℓ}}φ .

In particular, if φ is a parallel spinor, then the diagram

Γ ^{2} T^{}M Γ ^{2} T^{}M Γ Σ^{g}M T^{}M Γ Σ^{g}M T^{}M_{g,φ}_{g,φ} Δ_{E} D^{T^{}M} ^{2}

commutes.

Proof. We can locally write

g,φ h _{i,j} h_{ij}e_{j} φ e_{i}^{}

where e_{1}, … e_{n} is a local orthonormal frame with respect to g with e_{i} 0

at p. We define _{k}h_{ij} _{e_{k}}h e_{i}, e_{j} , _{k,l}^{2}h_{ij} _{e_{k},e_{l}}^{2}h e_{i}, e_{j} , and let

R_{e_{k},e_{l}}h_{ij}  _{k,l}^{2}h_{ij}  _{l,k}^{2}h_{ij} be the associated curvature. By using the Clifford

relations, we get on the domain of the frame:

D^{T^{}M} _{g,φ} h  _{i,j,k} _{k}h_{ij}e_{k} e_{j} φ h_{ij}e_{k} e_{j} _{e_{k}}φ e_{i}^{}

_{i,j,k} _{k}h_{ij}e_{k} e_{j} φ h_{ij}e_{j} e_{k} _{e_{k}}φ 2h_{ij}δ_{kj} _{e_{k}}φ e_{i}^{}

\begin{matrix}_{i,j,k} _{k}h_{ij}e_{k} e_{j} φ e_{i}^{} \\ \end{matrix}_{A} \begin{matrix}_{g,Dφ} h \\ \end{matrix}_{B} \begin{matrix}2 _{i,j} h_{ij} _{e_{j}}φ e_{i}^{} \\ \end{matrix}_{C}

Now we apply D^{T^{}M} to this equation and we use the notation ψ Dφ. Then we get from the previous equation at the point p

D^{T^{}M} B  D^{T^{}M} _{g,ψ} h

_{i,j,k} _{k}h_{ij}e_{k} e_{j} ψ e_{i}^{}_{g,Dψ} h 2 _{i,j} h_{ij} _{e_{j}}ψ e_{i}^{}.

Moreover,

D^{T^{}M} C  2 _{i,j,k} h_{ij}e_{k} _{e_{k},e_{j}}^{2}φ _{k}h_{ij}e_{k} _{e_{j}}φ e_{i}^{}

2 _{i,j,k} h_{ij}e_{k} _{e_{j},e_{k}}^{2}φ h_{ij}e_{k} R_{e_{k},e_{j}}φ _{k}h_{ij}e_{k} _{e_{j}}φ e_{i}^{}

2 _{i,j} h_{ij} _{e_{j}}ψ _{k} h_{ij}e_{k} R_{e_{k},e_{j}}φ _{k} _{k}h_{ij}e_{k} _{e_{j}}φ e_{i}^{}

and

D^{T^{}M} A \begin{matrix}_{i,j,k,l} _{lk}^{2}h_{ij}e_{l} e_{k} e_{j} φ e_{i}^{} \\  \end{matrix}_{E} \begin{matrix}_{i,j,k,l} _{k}h_{ij}e_{l} e_{k} e_{j} _{e_{l}}φ e_{i}^{} \\  \end{matrix}_{F} .

By using the relation e_{l} e_{k} e_{j} e_{k} e_{j} e_{l} 2e_{k}δ_{lj} 2δ_{kl}e_{j},

F _{i,j,k} _{k}h_{ij}e_{k} e_{j} ψ 2 _{k}h_{ij}e_{k} _{e_{j}}φ 2 _{k}h_{ij}e_{j} _{e_{k}}φ e_{i}^{}.

By adding up, a lot of terms cancel and we are left with

D^{T^{}M} ^{2} _{g,φ} h  E _{g,D^{2}φ} h 2 _{i,j,k} _{k}h_{ij}e_{j} _{e_{k}}φ e_{i}^{}

2 _{i,j,k} h_{ij}e_{k} R_{e_{j},e_{k}}φ e_{i}^{}

so it remains to consider the term E . We have

E  _{i,j,k} _{kk}^{2}h_{ij}e_{j} φ e_{i}^{}_{\begin{matrix}i,j,k,l \\ k l \end{matrix}} _{lk}^{2}h_{ij}e_{l} e_{k} e_{j} φ e_{i}^{}

_{i,j} ^{}h _{ij}e_{j} φ e_{i}^{}_{i,j,k,l} \frac{1}{2}R_{e_{k},e_{l}}h_{ij}e_{k} e_{l} e_{j} φ e_{i}^{}.

By the formula which expresses the curvature of T^{}M T^{}M in terms of the curvature of T M, we obtain R_{e_{k},e_{l}}h_{ij} _{r} R_{klir}h_{rj} R_{kljr}h_{ir} . Using the relation e_{k} e_{l} e_{j} e_{j} e_{k} e_{l} 2e_{k}δ_{jl} 2e_{l}δ_{jk}, we get

\frac{1}{2} _{i,j,k,l,r}R_{klir}h_{rj}e_{k} e_{l} e_{j}φ e_{i}^{}

\frac{1}{2} _{i,j,k,l,r} R_{klir}h_{rj}e_{j} e_{k} e_{l} φ e_{i}^{}

_{i,k,l,r} R_{klir}h_{rl}e_{k} R_{klir}h_{rk}e_{l} φ e_{i}^{}

\frac{1}{2} _{i,j,k,l,r} R_{klir}h_{rj}e_{j} e_{k} e_{l} φ e_{i}^{}2 _{i,j} R̊h _{ij}e_{j} φ e_{i}^{}

and

\frac{1}{2} _{i,j,k,l,r}R_{kljr}h_{ir}e_{k} e_{l} e_{j} φ e_{i}^{}

\frac{1}{2} _{i,j,k,l,r} R_{kljr}h_{ir}e_{j} e_{k} e_{l} φ e_{i}^{}

_{i,k,l,r} R_{kllr}h_{ir}e_{k} R_{klkr}h_{ir}e_{l} φ e_{i}^{}

\frac{1}{2} _{i,j,k,l,r} R_{kljr}h_{ir}e_{j} e_{k} e_{l} φ e_{i}^{}2 _{i,j,r} Ric_{jr} h_{ir}e_{j} φ e_{i}^{}.

We get

D^{T^{}M} ^{2} _{g,φ} h  _{g,φ} \begin{matrix}^{}h 2R̊h \\  \end{matrix}_{Δ_{E}h}  \frac{1}{2} _{i,j,k,l,r} R_{klir}h_{rj}e_{j} e_{k} e_{l} φ e_{i}^{}

\frac{1}{2} _{i,j,k,l,r} R_{kljr}h_{ir}e_{j} e_{k} e_{l} φ e_{i}^{}

2 _{i,j,r} Ric_{jr} h_{ir}e_{j} φ e_{i}^{}_{g,D^{2}φ} h

2 _{i,j,k} _{k}h_{ij}e_{j} _{e_{k}}φ e_{i}^{}2 _{i,j,k} h_{ij}e_{k} R_{e_{j},e_{k}}φ e_{i}^{}.

From the standard identities

R_{e_{i},e_{j}}φ \frac{1}{4} _{k,l} R_{ijkl}e_{k} e_{l} φ, _{i} e_{i} R_{e_{i},e_{j}}φ \frac{1}{2} Ric e_{j} φ,

we deduce

\frac{1}{2} _{i,j,k,l,r} R_{klir}h_{rj}e_{j} e_{k} e_{l} φ e_{i}^{}2 _{i,j,r} h_{rj}e_{j} R_{e_{i},e_{r}}φ e_{i}^{}
\frac{1}{2} _{i,j,k,l,r} R_{kljr}h_{ir}e_{j} e_{k} e_{l} φ e_{i}^{}2 _{i,j,r} h_{ri}e_{j} R_{e_{j},e_{r}}φ e_{i}^{}

_{i,j,r} h_{ri} Ric_{jr} e_{j} φ e_{i}^{}

2 _{i,j,k} h_{ij}e_{k} R_{e_{j},e_{k}}φ e_{i}^{}_{i,j,k} h_{ij} Ric_{jk} e_{k} φ e_{i}^{}

which yields the final result.

4 The BBGM parallel transport preserves parallel spinors

Proposition 8. Let I ℝ be an interval, g_{s} _{s I} be a path of Ricci-flat metrics with a divergence free derivative, 0 I, and let φ_{0} be a parallel spinor on M, g_{0} . Let, for all s I, φ_{s} be a spinor on M, g_{s} such that \frac{d}{ds}φ_{s} 0. Then φ_{s} is a parallel spinor on M, g_{s} for all s I.

Here \frac{d}{ds}φ_{s} 0 should be understood as a derivation with respect to the BBGM connection.

In the proof of Proposition 8, we need a statement about continuous dependence of eigenvalues of the Dirac operator on the metric.

Theorem 9. There exists a family of functions λ_{j} ℝ, j ℤ such that

• λ_{j} g _{j ℤ} spec D_{g} for all g

• The family is nondecreasing, i.e. λ_{i} λ_{j} whenever i j.

• sup_{j ℤ} arsinh λ_{j} g arsinh λ_{j} g̃ C g g̃ _{C^{1}}

The theorem in the above version is proven in detail in [33, Main Theorem 2], but the statement we need was well-known long before.

To prove the proposition, we will at first prove the same statement under more restrictive assumptions.

Lemma 10. Let g_{s} _{s I} be a path of Ricci-flat metrics with a divergence free derivative, 0 I, and let φ_{0} be a parallel spinor on M, g_{0} . Assume that dim ker D^{T^{}M} is constant along I. Let φ_{s} be a spinor on M, g_{s} such that \frac{d}{ds}φ_{s} 0. Then φ_{s} is a parallel spinor on M, g_{s} for all s I.

Proof of Lemma 10. Without loss of generality we can assume that I is compact. Using Rayleigh-Ritz type arguments, explained e.g. in detail in [33], one can show the existence of continuous functions I s λ_{i} s ℝ, λ_{i} s λ_{i 1} s such that λ_{i} s _{i ℤ} are the eigenvalues (including the correct multiplicity) of D^{T^{}M}. If the kernel of D^{T^{}M} has constant dimension k, then we can assume λ_{0} s λ_{1} s λ_{k} s 0 λ_{k 1} s . Then for

γ min_{s I}min λ_{0} s , λ_{k 1} s 0

the operator D_{g_{s}}^{T^{}M} has no eigenvalue in γ, 0 0, γ . Because of this, there are bounded operators

D_{g_{s}}^{T^{}M} ^{1} L_{g_{s}}^{2} ker D_{g_{s}}^{T^{}M} ^{}L_{g_{s}}^{2} ker D_{g_{s}}^{T^{}M} ^{}(18)

that invert D_{g_{s}}^{T^{}M} on the orthogonal complement of the kernel and that are uniformly bounded by γ^{1}. By elliptic theory, see e.g. [26, Chapter III §5] we have isomorphisms

D_{g_{s}}^{T^{}M} H_{g_{s}}^{1} ker D_{g_{s}}^{T^{}M} ^{}L_{g_{s}}^{2} ker D_{g_{s}}^{T^{}M} ^{}.  (19)

Because of im D_{g_{s}}^{T^{}M} ker D_{g_{s}}^{T^{}M} ^{}, there is a constant C_{0} 0 such that for any s I and for any σ Γ Σ^{g_{s}}M T^{}M we have

D_{g_{s}}^{T^{}M}σ _{H_{gs}^{1}} C_{0} D_{g_{s}}^{T^{}M} ^{2} σ _{L_{gs}^{2}} .  (20)

Moreover, g_{s}^{}ker Δ_{E,g_{s}} by the facts collected in Subsection 2.1. We calculate using Lemma 7

D_{g_{s}}^{T^{}M} ^{2} _{g_{s},φ_{s}} g_{s}^{}2 _{ℓ} _{g_{s}, _{e_{ℓ}} φ_{s}}  _{e_{ℓ}}g_{s}^{}_{g_{s},D^{2}φ_{s}} g_{s}^{}

2 X _{ℓ} g_{s}^{}^{}e_{ℓ} R_{X,e_{ℓ}}φ_{s} ,

thus pointwise D_{g_{s}}^{T^{}M} ^{2} _{g_{s},φ_{s}} g_{s}^{}is a linear expression in φ_{s} and its first derivative. Thus

D_{g_{s}}^{T^{}M} ^{2} _{g_{s},φ_{s}} g_{s}^{}_{L^{2}} C φ_{s} _{H^{1}},

where C depends on sup_{s I} g_{s}^{}_{C^{1} g_{s}}. This implies, using Eq. (20):

D_{g_{s}}^{T^{}M} _{g_{s},φ_{s}} g_{s}^{}_{H^{1}} CC_{γ} φ_{s} _{H^{1}}.  (21)

We now differentiate ^{g_{s}}φ_{s} using formulae (16) and (17) and Lemma 6.

\frac{d}{ds} ^{g_{s}}φ_{s} κ̂_{g_{s},φ_{s}} g_{s}^{}^{g_{s}} \begin{matrix}φ_{s}^{} \\  \end{matrix}_{0}

\frac{1}{4}{ \atop }D_{g_{s}}^{T^{}M} _{g_{s},φ_{s}} g_{s}^{}φ_{s} \begin{matrix}div^{g_{s}} g_{s}^{} \\ \end{matrix}_{0} 2 _{g_{s}^{}^{#}}φ_{s} g_{s}^{}^{#} D^{g_{s}}φ_{s}{ \atop }

\frac{1}{4}D_{g_{s}}^{T^{}M} _{g_{s},φ_{s}} g_{s}^{}\frac{1}{2} _{g_{s}^{}^{#}}φ_{s} \frac{1}{4} g_{s}^{}^{#} D^{g_{s}}φ_{s}.

This implies, with Eq. 21:

\frac{d}{ds} ^{g_{s}}φ_{s} _{H^{1} g_{s}} C̃{ \atop }^{g_{s}}φ_{s} _{H^{1} g_{s}} \begin{matrix}D^{g_{s}}φ_{s} _{H^{1} g_{s}} \\  \end{matrix}_{\overline{n} ^{gs} φ_{s} _{H^{1} gs}} { \atop }C_{1} ^{g_{s}}φ_{s} _{H^{1} g_{s}}

where C̃ and C_{1} depend on sup_{s I} g_{s}^{}_{C^{1} g_{s}}.

We set ℓ s ^{g_{s}}φ_{s} _{H^{1} g_{s}}^{2}. In order to derive ℓ s one has to be aware that also the evaluating metric depends on s. The dependence on the metrics effects ℓ s in the metric contractions, the volume element and in the covariant derivatives used to define by σ _{H^{1} g}^{2} σ _{L^{2} g}^{2} ^{g}σ _{L^{2} g}^{2}. Due to compactness of I, these effects lead to a number C_{2} 0, constant in s, such that

ℓ^{}s 2 \frac{d}{ds} ^{g_{s}}φ_{s}, ^{g_{s}}φ_{s} _{H^{1} g_{s}} C_{2}ℓ s

and we get

ℓ^{}s 2 \frac{d}{ds} ^{g_{s}}φ_{s}, ^{g_{s}}φ_{s} _{H^{1} g_{s}} C_{2}ℓ s 2C_{1} C_{2} ℓ s .

As ℓ 0 0 this implies with Grönwall’s inequality that ℓ s vanishes for all s.

Remark 11. In the proof above the derivative \frac{d}{ds} ^{g_{s}}φ_{s} should be taken with some care. Here we derive an s-dependent family ^{g_{s}}φ_{s} Γ T^{}M Σ^{g_{s}}M with respect to the metric variation given by s. This is the BBGM-derivative in the spinorial part. On the cotangential part, one could also use a BBGM-kind of derivative, but this is not what was used. On the cotangential part, we simply used the derivative in the usual sense, i.e. the derivative of a curve in the vector space Γ T^{}M .

Proof of Proposition 8. In a first step we prove the Proposition for analytic families g_{s}, s I of Ricci-flat metrics, in other words we assume that the map I ^{2} T^{}M defined by g_{s} is analytic.

Then D_{g_{s}}^{T^{}M} is an analytic family of operators. This implies that the eigenvalues λ_{i} s of D_{g_{s}}^{T^{}M} can be numbered such that s λ_{i} s is analytic in s [21, Appendix A]. Due to Theorem 9, only finitely many λ_{i} s have zero sets on a compact interval. Let μ min dim ker D_{g_{s}}^{T^{}M}, and J s I dim ker D_{g_{s}}^{T^{}M} μ which is a closed discrete subset due to analyticity. (We conjecture that J , but we were unable to prove it.) Let I_{0} be a connected component of I J. The previous lemma states that the Bourguignon-Gauduchon parallel transport along s g_{s} maps parallel spinor φ Γ Σ^{g_{s}}M to a parallel spinor P_{s,r} φ Γ Σ^{g_{r}}M for any s, r I_{0}. Let r̄ ∂I_{0}. Then by continuity,

P_{s,r̄} φ lim_{r r̄}P_{s,r} φ

is a parallel section of Γ Σ^{g_{r̄}}M . We use the fact that the dimension of the space of parallel spinors is locally constant (see [3]), thus for any s I_{0} and and r Ī_{0} (including r ∂I_{0}) the monomorphism P_{s,r} from parallel spinors in Γ Σ^{g_{s}}M to parallel spinors in Γ Σ^{g_{r}}M is an isomorphism. Thus BBGM parallel transport perserves parallel spinors along Ī_{0}, and by an induction argument (using that J is finite in compact intervals) this also holds along I. The proposition is thus proven for analytic families s g_{s} and thus also for piecewise analytic families s g_{s}.

Now, as the Riemannian metrics form an open cone in the vector space of bilinear forms, an arbitrary smooth family s g_{s} can be approximated by piecewise analytic paths in the C^{1}-norm. We claim that the BBGM parallel transport is continuous in this limit. This can be seen most easily in the universal spinor bundle formulation: There, for SM being the bundle of symmetric positive definite bilinear forms, a Clifford bundle π^{univ} ΣM SM is constructed such that, for each g Γ SM , Σ^{g}M is isomorphic as a Clifford bundle to g^{}ΣM. Furthermore, the vector bundle π^{univ} ΣM SM carries a vertical (w.r.t. SM M) covariant derivative whose parallel transport P is linked to the BBGM parallel transport P̂ as follows: Let g_{}0, ℓ Γ SM , s g_{s} be a C^{1}-curve of Riemannian metrics on M. Let p M and let g _{p} s g_{s} p and σ π^{univ} ^{1} g_{0} p Σ_{p}^{g_{0}}M, then for any σ̃ Γ Σ^{g_{0}}M with σ̃ _{p} σ we have

P̂_{g_{}}σ̃ _{p} P_{g _{p}} σ .

The parallel transport along a curve with respect to a connection is given by a first order ordinary differential equation (ODE), satisfying the conditions of the theorem of Picard-Lindelöf. Using the universal spinor bundle formalism we argued that the BBGM-parallel transport is given by such an ODE and that its coefficient functions converge uniformly (i.e. in the C^{0}-norm) when a path of metrics converges in the C^{1}-norm to a limit path of metrics.^{2} Thus the theorem of Picard-Lindelöf, taking into account that both M and 0, ℓ are compact, implies that the BBGM parallel transport converges uniformly when we approximate a smooth path of metrics s g_{s} by piecewise analytic paths of metrics in the C^{1}-norm. Thus the BBGM parallel transport also preserves parallel spinors along the smooth family s g_{s}.

5 Construction of solutions to the constraint equations

In this section we want to use the BBGM connection to construct solutions of the constraint equation on suitable manifolds of the form M I, g_{s} ds^{2} where I is an interval and where M is an m n 1 -dimensional manifold.

Assume that g_{s} is a family of Ricci-flat metrics with divergence-free derivative and that for some s_{0} there is a non-trivial parallel spinor φ_{s_{0}} on M, g_{s_{0}} . By rescaling we can achieve that its norm is 1 in every point. We shift it in the s-direction parallely with the BBGM parallel transport. By Prop. 8 we obtain a family φ_{s} of g_{s}-parallel spinors of constant norm 1 on M, for every s I. This yields a fiberwise parallel section of Σ_{M}^{#}N. Recall ν \frac{∂}{∂s}.

In the following we assume that f I ℝ is a given smooth function and choose s_{0} I. We define F s exp \frac{1}{2} _{s_{0}}^{s} f σ dσ , i.e. \frac{1}{2}f s F^{}s F s .

Case n is odd, i.e. m dim M even. By possibly changing the orientation and using Proposition 25 in Appendix C we can assume that all φ_{s} Γ Σ^{g_{s}}M have the positive parity for the splitting given by the volume element of M, i.e. we can assume ω_{ℂ}^{M} φ_{s} φ_{s}. For every x, s N M I we use the map J from Section 2.3 to get identifications Σ_{x}M Σ_{M}_{x,s}N Σ_{x,s}N. Again the Clifford multiplication for M, g_{s} will be denoted by , and for the one for N, h , h g_{s} ds^{2} we will use . From Equation (10) we get ν φ iφ. Using Lemma 4 we obtain

_{ν}^{N}φ  0,
_{X}^{N}φ  \frac{1}{2}W X ν φ \frac{i}{2}W X φ.

We set

Ψ x, s F s φ_{s} x .  (22)

Assuming that X is tangent to M, we obtain

ν Ψ  iΨ,
_{ν}^{N}Ψ  \frac{F^{}s}{F s}Ψ \frac{1}{2}f s Ψ \frac{i}{2}f s ν Ψ,
_{X}^{N}Ψ  \frac{1}{2}W X ν Ψ \frac{i}{2}W X Ψ.

g_{}max_{s 0,ℓ} g_{s} _{C^{0} M} \frac{d}{ds} g_{s} _{C^{0} M} .

2To be precise: we only need control of the norm

Thus Ψ is an imaginary \overline{W} -Killing spinor with \overline{W} W fν^{b} ν, i.e. it satisfies (3).

The first relation and the defining equation (22) also imply i ν Ψ, Ψ F^{2}. For X tangent to M, the real part of X Ψ, Ψ vanishes and

Im X Ψ, Ψ Re X iΨ, Ψ Re X ν Ψ, Ψ ^{}Re Ψ, X ν Ψ ,

where in we used the skew-symmetry of Clifford multiplication with vectors twice and X ν ν X. On the other hand one Re X ν Ψ, Φ Re Φ, X ν Ψ , which in particular holds for Φ Ψ. We thus get Im X Ψ, Ψ 0.

Thus i X Ψ, Ψ F^{2}h ν, X for every vector X and consequently, the Dirac current U_{Ψ} of Ψ, defined by h U_{Ψ}, X i X Ψ, Ψ for every vector X (see (1)) is U_{Ψ} F^{2}ν. We obtain

U_{Ψ} Ψ iF^{2}Ψ,

which is the constraint equation (4) for u_{Ψ} \overline{h U_{Ψ}, U_{Ψ}} F^{2}. Note that here we used (5) as definition for u_{ψ}. We thus have obtained solutions of the constraint equations.

Case n even, i.e. m dim M odd. We use the map J^{# ,} defined in Section 2.3 to view Σ_{x}M Σ_{M x,s}N as a subbundle of Σ_{x,s}N. In particular, x, s φ_{s} x then yields a section of constant length 1 of the bundle ΣN N.

Because of equations (12) and (13) we have ν φ iφ. Using Lemma 4 we obtain

_{ν}^{N}φ  0
_{X}^{N}φ  \frac{1}{2}W X ν φ \frac{i}{2}W X φ

We set

Ψ x, s F s φ_{s} x .  (23)

Then for X tangent to M,

ν Ψ  iΨ,
_{ν}^{N}Ψ  \frac{F^{}s}{F s}Ψ \frac{1}{2}f s Ψ \frac{i}{2}f s ν Ψ,
_{X}^{N}Ψ  \frac{1}{2}W X ν Ψ \frac{i}{2}W X Ψ.

Thus, Ψ is an imaginary \overline{W} -Killing spinor with \overline{W} W fν^{b} ν, i.e. it satisfies (3). With the same arguments as in the other case, we can prove that (1), (2), (4), and (5) are satisfied, i.e. we have found a solution to the constraint equations.

Example 12. Let M T^{n 1}, g be a flat metric on M and φ a parallel spinor on it.

• Let g_{s} g with s I. The BBGM parallel transport leaves φ invariant and we obtain initial data to the Cauchy problem on the metric g_{s} ds^{2} on either T^{n 1} ℝ or T^{n 1} S^{1} T^{n}. The Minkowski metric together with a parallel spinor (or in the S^{1} case a ℤ-quotient of it) is then a solution of the associated Cauchy problem.

• Let g_{s} e^{2s}g with s ℝ. For the submanifolds M s N we have W id _{T M} for alle s ℝ. We take the function f s 1, i.e. F s exp s 2 and \overline{W} id _{T N}. The metric we obtain is now the hyperbolic metric h e^{2s}g ds^{2} on T^{n 1} ℝ together with an imaginary Killing spinor with Killing constant i 2. The Lorentzian cone T^{n 1} ℝ ℝ_{0}, r^{2}h dr^{2} , where r ℝ_{0}, together with a parallel spinor solves solves the associated Cauchy problem. Note that this cone is the quotient by a ℤ^{n 1}-action of I_{}0 ℝ^{n,1}, defined as the set of all future-oriented time-like vectors in the n 1 -dimensional Minkowski space ℝ^{n,1}.

Remark 13. In this example we have seen two different ways of reconstructing Lorentzian Ricci-flat metrics on quotients of subsets of Minkowski space together with a parallel spinor. However, our construction allows many more interesting examples.

Example 14.

• Any imaginary Killing spinor on a complete, connected Riemannian spin manifold arises this way. This was proven by Baum, Friedrich, Grunewald and Kath in [7, Chap. 7], more precisely in Theorem 1 on page 160 and Cor. 1 on page 167 [7, Chap. 7].

• This was generalized by Rademacher [34], see also [18, Theorem A.4.5]. Rademacher proved that any generalized imaginary Killing spinor with W α id, α C^{}N, ℝ arises by our construction.

• Our construction generalizes previous constructions of imaginary Killing spinors, as e.g. [19, Prop. 4.6 and Cor. 4.7].

6 From curves in the moduli space to initial data sets

Let Diff_{0} M be the identity component of the diffeomorphism group of M, acting on the space _{}M of structured Ricci-flat metrics by pullback (as usual, a semi-Riemannian manifold M, g is called structured iff its semi-Riemannian universal covering ˜M, g̃ admits a parallel spinor). Furthermore, let od _{}M _{}M Diff_{0} M be the associated premoduli space. It was shown in [3] (using previous work about the special case of simply connected manifolds with irreducible holonomy) that od _{}M “naturally” carries the structure of a finite-dimensional smooth manifold. The smooth structure on od _{}M can be described by the following properties

• od _{}M carries the quotient topology, if we equip _{}M and Diff_{0} M with the standard Fréchet topology,

• any smooth family N _{}M , y g_{y} yields a smooth map N od _{}M , y g_{y} ,

• in the case N a, b one has \frac{d}{ds} _{s s_{0}} g_{s} 0 if and only if \frac{d}{ds} _{s s_{0}} g_{s} is tangent to the Diff_{0} M -orbit through g_{s_{0}}.

If I is an interval, then any smooth curve I od _{}M can be written as s g_{s} with a family of metrics div^{g_{s_{0}}} \frac{d}{ds} _{s s_{0}} g_{s} 0, see Appendix D.

The main results of this article is a procedure to construct initial data for the constraint equations for Lorentzian metrics with parallel spinors.

Main Construction 15 (Initial data on an open manifold). Let I be an interval. For any smooth curve I od _{}M s g_{s} and every positive smooth function F I ℝ_{0} we obtain a solution of the initial data equations (3) and (4) on M I with norm F t at any x, t M I.

Here the metric on M I is given by g_{s} ds^{2} provided that g_{s} is chosen such that div^{g_{s_{0}}} \frac{d}{ds} _{s s_{0}} g_{s} 0 and the spinor is given by (22) resp. (23). In particular the spinor can be normalized such that it has norm F t at any x, t M I. As derived in the preceding section, (3) are then satisfied, as well as (1), (2), (4), and (5).

The situation is slightly more complicated if we want to obtain solutions of the constraint equations on a closed manifold. We start with a closed curve S^{1} ℝ Lℤ od _{}M . If we identify ℝ Lℤ with 0, L 0 L, then every curve ℝ Lℤ od _{}M can be written as 0, L od _{}M , s g_{s} with div \frac{d}{ds}g_{s} 0, but in general we will have g_{0} g_{L} although M, g_{0} and M, g_{L} are isometric with respect to an isometry ζ Diff_{0} M . We then glue M 0 , g_{0} with M L , g_{L} isometrically using the diffeomorphism ζ Diff_{0} M . This yields a closed Riemannian manifold N, h diffeomorphic to M S^{1}. In order to equip it with a spin structure we have to lift dζ^{n} P_{SO} M, g_{0} P_{SO} M, g_{L} to a map between the corresponding spin structures ζ_{#} P_{Spin} M, g_{0} P_{Spin} M, g_{L} . This yields a spin structure and a spinor bundle on N. Let F S^{1} ℝ_{0} be given. For any parallel spinor φ_{0} on M, g_{0} equation (22) resp. equation (23) yields a generalized imaginary Killing spinor Ψ on M 0, L as in Main Construction 15. The gluing described above allows to view φ_{L} Ψ_{M L} as a parallel spinor on M, g_{0} . However, in general φ_{L} will differ from φ_{0}. Let Γ_{}Σ_{g_{0}}M denote the space of parallel spinors on M, g_{0} . Then φ_{0} φ_{L} yields a unitary map P Γ_{}Σ_{g_{0}}M Γ_{}Σ_{g_{0}}M . The map P does neither depend on F nor on the parametrization of the curve s g_{s} . We say that M, g_{s}, φ_{0} satisfies the fitting condition if P φ_{0} φ_{0} for a suitable choice of spin structure on N. The fitting condition is always satisfied in the following cases:

(1) M, g_{0} is a 7-dimensional manifold with holonomy G_{2}

(2) M, g_{0} is an 8-dimensional manifold with holonomy Spin 7

(3) M, g_{0} is a Riemannian product of manifolds of that kind and of at most one factor diffeomorphic to S^{1}.

(4) finite quotients of such manifolds

As this statement is not within the core of this article, we only sketch the proof. In the first two cases the spinor bundle Σ_{g}M is the complexification of the real spinor bundle Σ_{g}^{ℝ}M, and thus Γ_{}Σ_{g}M Γ_{}Σ_{g}^{ℝ}M ℂ. The real spinor representations of G_{2} and Spin 7 on Σ_{n}^{ℝ} have a 1-dimensional invariant subrepresentation, thus dim_{ℝ} Γ_{}Σ_{g}^{ℝ}M 1. Thus P is either id or id, and the -sign can be achieved by a suitable choice of the lift ζ_{#}. On S^{1} it follows from a direct calculation.

The map P behaves “well” under taking products and finite quotients, thus the other two statements follow as well.

For manifolds with a least one factor of holonomy SU k or Sp k , or also for tori of dimension 1, however, we expect that generically the fitting condition does not hold. In this case, we expect that the space of closed paths s g_{s} for wich P has finite order (in the sense ℓ ℕ P^{ℓ} id) is dense in the space of all closed paths s g_{s} with respect to the C^{}-topology. This is in fact a consequence of work in progress by Bernd Ammann, Klaus Kröncke and Hartmut Weiß.

Then passing to an ℓ-fold cover of N obtained from running along the path s g_{s} not just once, but ℓ times, we obtain a solution of the constraint equation on M S^{1}.

Main Construction 16 (Initial data on a closed manifold). Let L 0. Let ℝ Lℤ od _{}M , s g_{s} be a smooth path and let φ_{0} Γ_{}Σ^{g_{0}}M be given such that M, g_{s}, φ_{0} satisfies the fitting condition. Then for any function F S^{1} ℝ_{0} we obtain a solution of the initial data equations (3) and (4) on M S^{1}.

It also seems interesting to us to allow a slight generalization of our initial problem, by considering spin^{c} spinors with a flat associated line bundle instead of spinors in the usual sense. Assume that ϑ ℂ has norm 1. Identifying v, t ℂ ℝ with ϑv, t L yields a complex line bundle L_{ϑ} over S^{1} ℝ Lℤ. On L_{ϑ} we choose the connection such that local sections with constant v are parallel. By pull back we obtain complex line bundles with flat, metric connections on N M S^{1} and on \overline{N} N ϵ, ϵ . These line bundles will also be denoted by L_{ϑ}. The bundle Σ_{h}N L_{ϑ} resp. Σ_{h̄}\overline{N} L_{ϑ} is then a spin^{c}-spinor bundle with flat associated bundle L_{ϑ}. The objects tensored by L_{ϑ} will be called L_{ϑ}-twisted. All the results of this article immediately generalize to L_{ϑ}-twisted spinors. We ask for L_{ϑ}-twisted parallel spinors, i.e. parallel sections of Σ_{h̄}\overline{N} L_{ϑ} instead of parallel spinors in the usual sense. This leads to L_{ϑ}-twisted constraint equations, and the L_{ϑ}-twisted Cauchy problem can be solved the same way as the untwisted.

Let P U Γ_{}Σ_{g_{0}}M be as above. As P is unitary, there is a basis of Γ_{}Σ_{g_{0}}M consisting of eigenvectors of P for complex eigenvalues of norm 1.

Main Construction 17 (Spin^{c}-version). Let L 0. Let ℝ Lℤ od _{}M , s g_{s} be a smooth path. Let φ_{0} Γ_{}Σ_{g_{0}}M be an eigenvalue of P to the eigenvalue ϑ. Then for any function F S^{1} ℝ_{0} we obtain a solution of the L_{ϑ}-twisted version of the constraint equations (3) and (4) on M S^{1}.

A Independence of the constraint equations

In this appendix we want to show that the constraint equations (1)–(5) presented in the introduction are not independent equations. We will show that all of them follow from (3) and a rewritten version of (4). In particular, we will see that for a generalized imaginary Killing spinor φ equation (4) implies (2), unless the vector field U_{φ} vanishes everywhere. In the introduction (1)–(5) are a mixture of definitions and relations. Let us rewrite them in a form which is more suitable to clarify their dependences.

We assume that N, h is a connected Riemannian spin manifold. Let ΣN N be the associated spinor bundle. Compared to the introduction we slightly simplify our notation: we write here for the Clifford multiplication instead of writing which was used in the introduction in order to distinguish it from other Clifford multiplications.

In the following we have a spinor φ, i.e. a smooth section φ Γ ΣN , a real-valued smooth function u C^{}N , a vector field U Γ T N and an endomorphism W Γ T N . Be aware that we simplify again the notation, by writing W for the endomorphism which was called \overline{W} in the introduction. The equations (1)–(5) turn into

h U, X i X φ, φ ,  X T N, ,  (24)
iφ p V φ V T_{p}N ,  p N,  (25)
_{X}^{N}φ \frac{i}{2}W X φ,  X T N,  (26)
U φ iu φ,  (27)
u^{2} h U, U .  (28)

We want to discuss the independence of the constraint equations (24) to (28). We start with some elementary lemmata.

Lemma 18. If (27) is satisfied for U, u and φ defined in some p N. Then we have (in this p N):

φ 0 or h U, U u^{2}

Proof.

h U, U φ, φ U U φ, φ U φ, U φ iu φ, iu φ u^{2} φ, φ .

Lemma 19. Let p N and let (27) be satisfied for U T_{p}N, u ℝ and φ Σ_{g}N_{p}. We assume that U_{φ} is defined by (1), i.e. (24) holds for U_{φ} instead of U at the point p. Then U and U_{φ} are linearly dependent.

In that sense (27) implies (24) up to a constant. Obviously for globally defined U, u and φ, the proportionality factor does not have to be constant. We obtain λ_{1}U λ_{2}U_{φ} for some nowhere vanishing function λ M ℝ^{2}.

Proof. W.l.o.g. U 0, φ 0 at p N. By Lemma 18 it follows that u p 0. We calculate for X U:

0  2h X, U φ, φ X U φ, φ U X φ, φ
X U φ, φ φ, X U φ 2 Re X U φ, φ .

Furthermore

X U φ, φ iu X φ, φ iu X φ, φ u h U_{φ}, X .

This implies h U_{φ}, X 0.

Lemma 20 (In [9, Lemma 5 in Sec. 5.2]). Assume U and φ satisfy (24). Then (25) is equivalent to h U, U φ ^{4}.

Proof. We prove the statement in each p N, so we consider U T_{p}N and φ Σ_{p}N. W.l.o.g. φ 0. If e_{j} is an orthonormal basis of T_{p}N, then ρ_{j} with ρ_{j} \frac{1}{φ}e_{j} φ is an orthonormal basis of E_{φ} V φ V T_{p}N . We write

iφ W φ ψ

for some W T_{p}N and ψ V φ V T_{p}N . Thus

iφ, ρ_{j} \frac{i}{φ} φ, e_{j} φ \frac{1}{φ}h U, e_{j} .

We conclude

h U, U _{j 1}^{n} h U, e_{j} ^{2} φ ^{2} _{j 1}^{n} iφ, ρ_{j} ^{2} φ ^{2} iφ ^{2} ψ ^{2} .

This implies that h U, U φ ^{4} if and only if ψ 0.

Now let W End T N , not necessarily symmetric. Recall that our notation is slightly simplified if compared to the introduction: the \overline{W} of the introduction is W in this appendix.

Proposition 21 (Dichotomy Proposition). Let N, h be a connected Riemannian manifold and let W Γ End T N be a field of endomorphisms. We assume that U U_{φ} for φ satisfy (24), (26) and (27) for some u C^{}N .

Then W^{T} U_{φ} grad φ, φ . If U_{φ} 0, then φ is constant. If U_{φ} 0, then U_{φ} and φ vanish nowhere and u U_{φ} φ, φ φ ^{2}.

Here W^{T} denotes the endomorphism in End T N adjoint to W.

Let us compare this propositition to a similar statement by H. Baum and Th. Leistner. In the case that W is symmetric, it yields a criterion implying (27).

Lemma 22 ([9, Lemma 5 in Sec. 5.2]). Let N, h be a connected Riemannian manifold with a non-zero spinor field φ Γ ΣN and a field of symmetric endomorphisms W Γ End T N satisfying (26), and let U U_{φ} be defined by (24). Then we have U_{φ} φ ^{2}W. Furthermore q_{φ} φ ^{4} h U_{φ}, U_{φ} is nonnegative and constant. Moreover, if we define ψ as in the proof of Lemma 20, then q_{φ} φ ^{2} ψ ^{2}. If q_{φ} 0, then (27) holds for u U_{φ} φ ^{2}.

Note that in the case of an imaginary Killing spinor, i.e. W μ id, then q_{φ} is related to the constant Q_{φ} defined in [7] for any twistor spinor by Q_{φ} n^{2}μ^{2}q_{φ}. Note that imaginary Killing spinors are both twistor spinors and generalized imaginary Killing spinors, but there are generalized imaginary Killing spinors, which are not twistor spinors and vice versa. According to [7, Chap. 7] an imaginary Killing spinor is of type I, if and only if Q_{φ} 0. Otherwise it is of type II. Any complete Riemannian manifold carrying a type II imaginary Killing spinor (with μ 0) is homothetic to the hyperbolic space [7, Sec. 7.2]. If it is of type I, then it arise from a warped product construction as in our Section 5, see [7, Sec. 7.3].

Proof of Proposition 21. First we compute

X φ, φ  _{X}φ, φ φ, _{X}φ  \frac{i}{2} W X φ, φ φ, W X φ
i W X φ, φ h U_{φ}, W X
h W^{T} U_{φ} , X .

With X φ, φ h grad φ, φ , X we obtain grad φ, φ W^{T} U_{φ} . Obviously this implies that φ is constant in the case U_{φ} 0.

We now consider the case U_{φ} 0. Note that Equation (26) is equivalent to saying that φ is a parallel section for the connection \overline{}_{X}φ _{X}^{N}φ \frac{i}{2}W X φ. This implies: If we have some p M with φ p 0, then φ has to vanish on all of N, and thus U_{φ} 0, the case already solved. So let us assume φ p 0 for all p M. Note that (27) implies

U_{φ} \frac{U_{φ} φ}{φ} \frac{iuφ}{φ} u .

We calculate

u^{2} φ, φ iu φ, iu φ  U_{φ} φ, U_{φ} φ
h U_{φ}, U_{φ} φ, φ
i U_{φ} φ, φ φ, φ
i iu φ, φ φ, φ
u φ, φ ^{2},

i.e. at each point p N we have u p 0 or u p φ p , φ p φ p ^{2}. The sets p N u p 0 and p N u p φ p ^{2} are closed and disjoint, thus the connectedness of N implies that either u 0 or u φ, φ . In the case u 0 we obtain U_{φ} 0, and we are again back in the case already solved. So we conclude u φ, φ φ ^{2} 0 and thus u U_{φ} . So everything is proven.

Now we discuss our main case of interest, i.e. that φ is a generalized imaginary Killing spinor which is by definition a solution of _{X}φ \frac{i}{2}W X φ X T N with W End T N symmetric. We assume N to be connected and φ 0 which implies as we have seen that φ vanishes nowhere. According to Lemma 22 we obtain the equation U_{φ} φ ^{2}W and the fact that q_{φ} φ ^{4} U_{φ} ^{2} is a non-negative constant. In the case q_{φ} 0 (denoted by “type I” in [9, Sec. 5.2]) we know further that (27) holds for u U_{φ} .

Corollary 23. Let N, h be a connected Riemannian spin manifold. Let φ be a generalized imaginary Killing spinor, i.e. a solution of (26) for a field of symmetric endomorphisms W. Let again U U_{φ} be defined by (24). We assume W 0, thus U_{φ} 0 and hence U_{φ} 0. Then the following are equivalent:

(a) φ is of type I,

(b) φ satisfies U_{φ} φ ^{2},

(c) φ satisfies equation (2) (or equivalently (25)), (d) Condition (27) holds for u U_{φ} .

Proof.

(a) (b) holds by definition of “type I”.

(b) (c) is stated in Lemma 20.

(b) (d) is part of the statement of Lemma 22.

(d) (b) is part of the statement of Proposition 21.

We have seen, in particular, that for a generalized imaginary Killing spinor φ equation (27) implies (25) for U U_{φ} defined by (24), unless the vector field U_{φ} and the endomorphism W vanish everywhere.

B More on Hypersurfaces

In this appendix we prove Lemma 4, i.e. formula (14).

It is known since long that one cannot restrict spinors to a hypersurface in a way preserving the connection. The difference of the connections depends on the second fundamental form or equivalently the Weingarten map. This effect is in some applications very helpful, e.g. in the case of surfaces in Euclidean space, where it leads to the spinorial version of the Weierstrass representation, see [17] for a good presentation or see [25] for an earlier, up to branching point aspects complete, but less conceptual publication, based on [24]. How to restrict spinors to hypersurfaces and the effect on the connection was already discussed in mathematical physics in the Riemannian [37, 38] and Lorentzian [40] context, and in spectral theory [5].

As different convention are used in the literature and as we follow, similar to [1, Prop. 5.3.1], another convention than the well-written exposition [6] we want to give a detailed proof of Lemma 4 in this appendix.

As in Subsection 2.3 and Appendix B we assume that N, h is an ndimensional Riemannian spin manifold, N M a, b , h g_{s} ds^{2} for a family of metrics g_{s} , s a, b on M. Let ^{M,g_{s}} be the Levi-Civity connection of M, g_{s} and ^{N} the one of N, h . We write ν ∂ ∂s for the unit normal vector field of M s in N. For X T M T M s T N, and Y Γ T M we have

_{X}^{N}Y _{X}^{M,g_{s}}Y II X, Y , g_{s} W X , Y ν II X, Y ,

and we have, see e.g. [6, (4.1)] g_{s} W X , Y \frac{1}{2} d ds g_{s}. Note that the second fundamental form II has values in the normal bundle. Again, we use for the Clifford multipication on N.

For concrete calculations we choose an open subset U M a, b N and for every x, s U we choose a positively oriented g_{s}-orthonormal basis q_{M} e_{1}, …, e_{n 1} of T_{x}M, smoothly depending on x and s. Thus q_{M} is a local section of P_{SO,M} N . In other words: q_{M} is a frame of the vertical bundle of N a, b . Furthermore q_{N} e_{1}, …, e_{n} , e_{n} ν is a frame for N, h , i.e. a local section of P_{SO} N . We define the associated Christoffel symbols by

_{e_{i}}^{M,g_{s}}e_{j} _{k 1}^{n 1} Γ_{ij}^{M k}e_{k}, _{e_{i}}^{N}e_{j} _{k 1}^{n} Γ_{ij}^{N k}e_{k}.

Now let q̃_{M} resp. q̃_{N} be a spinorial lift of q_{M} resp. q_{N}, i.e. a local section of P_{Spin,M} N resp. P_{Spin} N , such that postcomposing q̃_{M} resp. q̃_{N} with P_{Spin,M} N P_{SO,M} N resp. P_{Spin} N P_{SO} N yields q_{M} resp. q_{N}. On U we can write a spinor Φ, i.e. a section of Φ Γ Σ^{#}N as Φ q̃_{N}, σ , with σ U Σ_{n}^{#}. We also may view as Σ^{#}N as an associated bundle to P_{SO,M} N , and with proper identifications we get q̃_{N}, σ q̃_{M}, σ . The connection ^{N} defines the standard Levi-Civita connection on ΣN, again denoted by ^{N}. On the other hand, ^{M,g_{s}} defines a connection on ΣN _{M s} .

Proposition 24. For any i 1, …, n 1 and Φ Γ Σ^{#}N we have

_{e_{i}}^{N}Φ _{e_{i}}^{M,g_{s}}Φ \frac{1}{2} _{j 1}^{n 1} e_{j} II e_{i}, e_{j} Φ.

Proof. For any i, j, k 1, …, n 1 we have

Γ_{ij}^{N k} Γ_{ij}^{M k},  Γ_{ij}^{N n} e_{n} Γ_{in}^{N j} e_{n} II e_{i}, e_{j} ,  Γ_{in}^{N n} 0.

Let E_{1}, …, E_{n} denote the canonical basis of ℝ^{n}. Writing the connection in local coordinates we obtain for i 1, 2, …, n 1

_{e_{i}}^{N} q̃_{M}, σ  q̃_{M}, ∂_{e_{i}}σ \frac{1}{4} _{j,k 1}^{n} Γ_{ij}^{N k}E_{j} E_{k} σ

\begin{matrix}q̃_{M}, ∂_{e_{i}}σ \frac{1}{4} _{j,k 1}^{n 1} Γ_{ij}^{N k}E_{j} E_{k} σ \\  \end{matrix}_{_{e_{i}}^{M,gs} q̃_{M} ,σ} \begin{matrix}q̃_{M}, \frac{1}{4} _{j 1}^{n 1} Γ_{ij}^{N n}E_{j} E_{n} σ \\  \end{matrix}_{\frac{1}{4} _{j 1}^{n 1} e_{j} II e_{i},e_{j}  q̃_{M} ,σ}

\begin{matrix}q̃_{M}, \frac{1}{4} _{k 1}^{n 1} Γ_{in}^{N k}E_{n} E_{k} σ \\ \end{matrix}_{\frac{1}{4} _{j 1}^{n 1} e_{j} II e_{i},e_{j} q̃_{M} ,σ}

_{e_{i}}^{M,g_{s}} q̃_{M}, σ \frac{1}{2} _{j 1}^{n 1} e_{j} II e_{i}, e_{j}  q̃_{M}, σ .

Obviously we have the following calculation in the Clifford algebra

_{j 1}^{n 1} e_{j} II e_{i}, e_{j} _{j 1}^{n 1} e_{j} g_{s} W e_{i} , e_{j} ν W e_{i} ν,

and thus we obtain for X T M and Φ Γ Σ^{#}N the equation

_{X}^{N}Φ _{X}^{M,g_{s}}Φ \frac{1}{2}W X ν Φ.

As the maps J^{# ,} in Lemma 4 are constructed from an algebraic map using the associated bundle construction, they are ^{M,g_{s}}-parallel, i.e.

_{X}^{M,g_{s}}J^{# ,} φ J^{# ,} _{X}^{M,g_{s}}φ

for all X T M. Now Lemma 4 follows immediately by setting Φ J^{# ,} φ .

C Change of orientation

In this appendix we will prove the following proposition.

Proposition 25. Let M, g be a Riemannian manifold with an orientation , m dim M. Let P_{Spin} M, g, P_{SO} M, g, be a spin structure on M, g, with associated spinor bundle Σ M, g, P_{Spin} M, g, _{σ_{m}}Σ_{m} resp. Σ^{#} M, g, P_{Spin} M, g, _{σ_{m}} Σ_{m}^{#}. Then there is a spin structure P_{Spin} M, g, P_{SO} M, g, on M, g, with the following properties: Case m even: let Σ M, g, P_{Spin} M, g, _{σ_{m}} Σ_{m} be the associated spinor bundle, then there is a parallel isometric, complex linear bundle isomorphism Ψ Σ M, g, Σ M, g, , commuting with Clifford multiplication. This map Ψ maps Σ M, g, ^{}to Σ M, g, ^{}and Σ M, g, ^{}to Σ M, g, ^{}. Case m odd: let Σ^{#} M, g, P_{Spin} M, g, _{σ̃_{m}} Σ_{m}^{#} be one of the associated spinor bundles. Then there is a parallel isometric, complex linear bundle isomorphism Ψ Σ M, g, Σ^{#} M, g, , commuting with Clifford multiplication. The same statement holds if we exchange the role of Σ_{m} and Σ_{m}^{#} in the definitions of Σ M, g, .

We define a map ρ P_{SO} M, g, P_{SO} M, g, by ρ e_{1}, e_{2}, …, e_{m} : e_{1}, e_{2}, …, e_{m} for any -oriented orthonormal frame : e_{1}, …, e_{m} .

As a covering of smooth manifolds we define P_{Spin} M, g, P_{SO} M, g, as the pullback of P_{Spin} M, g, P_{SO} M, g, by the diffeomorphism ρ. Let ρ̃ P_{Spin} M, g, P_{Spin} M, g, be the diffeomorphism defined by the following pull-back square:

P_{Spin} M, g, P_{Spin} M, g, P_{SO} M, g, P_{SO} M, g, ρ̃ ρ

However in order to get an appropriate structure as Spin m -principal bundle on P_{Spin} M, g, , some care is necessary, as ρ is not SO m -equivariant. If we define J diag 1, 1, 1, 1, ...1 and abbreviate : e_{1}, …, e_{m} , then we have ρ J and thus ρ A ρ J^{1}AJ.

Conjugation with J J^{1} is a Lie group automorphism of SO m and lifts to Spin m Cl_{m}, as conjugation with E_{1} : 1, 0..., 0 in the Clifford algebra sense. For any P_{Spin} M, g, and B Spin m we define

B : ρ̃^{1} ρ̃ E_{1}B E_{1} .

This definition turns P_{Spin} M, g, into a Spin m -principal bundle and the map P_{Spin} M, g, P_{SO} M, g, is then Spin m SO m -equivariant.

Let as before σ Cl_{m} End Σ_{m}^{#} be an irreducible representation of the Clifford algebra, and let Σ^{#} M, g, : P_{Spin} M, g, _{σ} Σ_{m}^{#} with the obvious modifications for .

Lemma 26 (Lift to the spinor bundle). The map

P_{Spin} M, g,  Σ_{m}^{#}  , φ ρ̃ , σ E_{1} φ P_{Spin} M, g, Σ_{m}^{#}

is compatible with the equivalence relation given by σ. Thus it descends to a map

ρ_{#} Σ^{#} M, g,  P_{Spin} M, g,  _{σ} Σ_{m}^{#} Σ^{#} M, g, .

Proof. B, σ B^{1} φ is mapped to

ρ̃ B , σ E_{1} σ B^{1} φ ρ̃ E_{1}B E_{1} , σ E_{1}B E_{1} ^{1} σ E_{1} φ .

This pair is equivalent to ρ̃ , σ E_{1} φ which is the image of , φ .

Remark 27. For future publications it might be helpful here to briefly discuss, what happens if apply this change of orientation twice. Obviously, by replacing by we get another map ρ_{#} Σ^{#} M, g, Σ^{#} M, g, . It is easy to check that we then have ρ_{#} ^{2} id.

In the following sections of an associated vector bundle V P _{ρ} W — where P is a principal G-bundle and where W is a G-representation — are written as an equivalence class A, w of the pair A, w with respect to the action of ρ G GL W . Here A is a local section of P and w a locally defined function M W.

Lemma 28 (Compatibility with the Clifford action).

X ρ_{#} φ ρ_{#} X φ

for X T_{p}M, φ Σ_{p}M.

In particular, this lemma implies that although ρ_{#} yields an isomorphism between spinor bundles for different orientations, it does not yet have the properties that we request for Ψ.

Proof. We view T M as an associated bundle to P_{Spin} M, g, where either sign yields a possible description. In these descriptions , v and ρ̃ , Jv represent the same vector. Thus

, v ρ_{#} , φ ρ̃ , Jv ρ̃ , σ E_{1} φ
ρ̃ , σ Jv σ E_{1} φ
ρ̃ , σ E_{1} σ v φ
ρ_{#} , v , φ .

Here we used that Jv E_{1} E_{1} v in Cl_{m}.

Lemma 29. Let X T_{p}M, φ Γ Σ M, g, . Then

Xρ_{#} φ ρ_{#} _{X}φ .  (29)

Proof. The differential of ρ P_{SO} M, g, P_{SO} M, g, maps T P_{SO} M, g, to T P_{SO} M, g, . The connection 1-form ω SO M 𝔰𝔬 n then pulls back according to

ω dρ Y Jω Y J

for Y T_{}P_{SO} M, g, , a lift of X T M under the projection P_{SO} M, g, M. We lift this to a connection 1-form ω̃ P_{Spin} M, g, 𝔰𝔭𝔦𝔫 m Cl_{m} which thus transforms as

ω̃ dρ̃ Ỹ E_{1}ω Ỹ E_{1}

where Ỹ T P_{Spin} M, g, is a lift of Y. And this induces the relation (29).

Note that obviously ρ_{#} is isometric in each fiber.

Now finally in order to get a map Ψ as in the proposition, we will compose ρ_{#} with a further bundle isomorphism. We define σ̂ X : σ X for any X ℝ^{m}. Obviously, σ̂ satisfies the Clifford relations, and thus Σ_{m}^{#}, σ̂ describes a representation of Cl_{m} which is due to its dimension irreducible. The classification of such representations implies that this representation is isomorphic to either Σ_{m}, σ or (possibly in the case m odd) Σ_{m}^{#}, σ . If m is even, then we obtain a complex vector space isomorphism K End Σ_{m} satisfying

K σ X φ σ X K X  (30)

for all X ℝ^{m} and φ Σ_{m}. We can choose K to be isometric. Similarly, by checking the effect of Clifford multiplication by the volume element we obtain for m odd isometric complex isomorphisms K Σ_{m} Σ_{m}^{#} and K Σ_{m}^{#} Σ_{m} satisfying (30) for X ℝ^{m} and φ Σ_{m} resp. φ Σ_{m}^{#}. The associated map defined by K defines parallel, fiberwise isometric complex linear isomorphisms of vector bundles over id_{M}

K Σ M, g, Σ M, g,

for m even and

K Σ M, g,  Σ^{#} M, g, , K Σ^{#} M, g,  Σ M, g,

for m odd, satisfying (30) for X T_{p}M and φ Σ^{#} M, g, .

The composition Ψ K ρ_{#} now satisfies all properties requested in the proposition.

Remark 30. Let m be even. Recall that a spinor φ is called positive resp. negative if ω_{ℂ}φ φ resp. ω_{ℂ}φ φ. It is easy to check that ρ_{#} and Ψ map positive spinors to negative ones and vice versa, while K preserves positivity and negativity of spinors.

The proof of Proposition 25 is thus complete.

D Making paths of metrics divergence free

In this appendix we show the following well-known lemma.

Lemma 31. Let g_{s}, s 0, ℓ be a path of Riemannian metrics on a closed manifold M. We assume that the dimension of the space of Killing vector fields of M, g_{s} does not depend on s. Then there exists a family of diffeomorphisms φ_{s} M M, depending smoothly on s 0, ℓ , φ_{0} id_{M}, such that g̃_{s} φ_{s}^{}g_{s} satisfies for all s 0, ℓ :

div^{g̃_{s}} \frac{d}{ds}g̃_{s} 0.  (31)

Proof. We make the following ansatz. Let X_{s} Γ T M be a vector field smoothly depending on the parameter s 0, ℓ . Let φ_{s} be the flow generated by X_{s}, i.e.

\frac{d}{ds}φ_{s} x X_{s} _{φ_{s} x}.

For this we define h_{s} \frac{d}{ds}g_{s}. Then

\frac{d}{ds}g̃_{s} φ_{s}^{}_{X_{s}}g_{s} h_{s} ,

where _{X_{s}} denotes the Lie derivative in the direction of X_{s}. Thus (31) is equivalent to

div^{g_{s}} _{X_{s}}g_{s} h_{s} 0.

Now let div^{g_{s}} ^{}Ω^{1} M Γ T^{}M T^{}M be the adjoint of div^{g_{s}}. Then, see [12, Lemma 1.60],

Xg 2 div^{g} ^{}X^{b}

for all X Γ T M and Riemannian metrics g, where X^{b} g X, . We thus see that (31) is in fact equivalent to

2 div^{g_{s}} div^{g_{s}} ^{}α_{s} div^{g_{s}} h_{s}.  (32)

where we set α_{s} g_{s} X_{s}, . By calculation the principal symbol of P_{s} div^{g_{s}} div^{g_{s}} ^{}one sees that P_{s} is a self-adjoint elliptic operator, thus has discrete (non-negative) spectrum. We have ker P_{s} ker div^{g_{s}} ^{}. Thus again by [12, Lemma 1.60] the kernel of P_{s} is the space of all Killing vector fields of M, g_{s} . We furthermore have

im P_{s} ker P_{s} ^{}ker div^{g_{s}} ^{}^{}im div^{g_{s}} .

thus (32) has a unique solution α_{s} that is L^{2}-orthogonal to any Killing vector field. By assumption, the dimension of ker P_{s} is constant. Thus, the spaces im P_{s} form a smooth family of isomorphic vector spaces and we have a smooth family of isomorphisms P_{s} on im P_{s} . Thus, α_{s} and hence also X_{s} and φ_{s} depend smoothly on s. This solves the problem.

Note that we apply the above theorem to a family of Ricci-flat metrics. On closed Ricci-flat Riemannian manifolds every Killing vector field is parallel. Furthermore X is then parallel if and only if X^{b} is harmonic. Thus the dimension of the space of Killing vector fields is the first Betti-number and thus independent of s.

References

[1] B. Ammann; A Variational Problem in Conformal Spin Geometry, Habilitationsschrift, Universität Hamburg, 2003; URL

http://www.mathematik.uni-regensburg.de/ammann/preprints/habil/habil.ps.

[2] B. Ammann, J. Glöckle; Dominant energy condition and spinors on Lorentzian manifolds (2021); preprint, arXiv:2103.11032.

[3] B. Ammann, K. Kröncke, H. Weiss, F. Witt; Holonomy rigidity for Ricci-flat metrics; Math. Z. 291, 303–311 (2019); URL

http://dx.doi.org/10.1007/s00209-018-2084-3.

[4] B. Ammann, H. Weiss, F. Witt; A spinorial energy functional: critical points and gradient flow; Math. Ann. 365, 1559–1602 (2016); URL

http://dx.doi.org/10.1007/s00208-015-1315-8.

[5] C. Bär; Extrinsic bounds for eigenvalues of the Dirac operator; Ann. Global Anal. Geom. 16, 573–596 (1998).

[6] C. Bär, P. Gauduchon, A. Moroianu; Generalized Cylinders in Semi-Riemannian and Spin Geometry; Math. Z. 249, 545–580 (2005).

[7] H. Baum, T. Friedrich, R. Grunewald, I. Kath; Twistors and Killing spinors on Riemannian manifolds; Teubner-Texte zur Mathematik [Teubner Texts in Mathematics], vol. 124, B. G. Teubner Verlagsgesellschaft mbH, Stuttgart 1991; with German, French and Russian summaries.

[8] H. Baum, K. Lärz, T. Leistner; On the full holonomy group of Lorentzian manifolds; Math. Z. 277, 797–828 (2014); URL

http://dx.doi.org/10.1007/s00209-014-1279-5.

[9] H. Baum, T. Leistner; Lorentzian Geometry — Holonomy, Spinors, and Cauchy Problems; in V. Cortes, K. Kröncke, J. Louis (eds.), Geometric Flows and the Geometry of Space-time; Tutorials, Schools, and Workshops in the Mathematical Sciences, Birkhäuser, 2018; URL

https://www.springer.com/us/book/9783030011253.

[10] H. Baum, T. Leistner, A. Lischewski; Cauchy problems for Lorentzian manifolds with special holonomy; Differential Geom. Appl. 45, 43–66 (2016);

URL http://dx.doi.org/10.1016/j.difgeo.2015.11.007.

[11] L. Bérard-Bergery, A. Ikemakhen; On the holonomy of Lorentzian manifolds; in Differential geometry: geometry in mathematical physics and related topics (Los Angeles, CA, 1990); Proc. Sympos. Pure Math., vol. 54; 27–40, Amer. Math. Soc., Providence, RI 1993.

[12] A. L. Besse; Einstein Manifolds; no. 10 in Ergebnisse der Mathematik und ihrer Grenzgebiete, 3. Folge, Springer 1987.

[13] C. Boubel; On the holonomy of Lorentzian metrics; Ann. Fac. Sci. Toulouse Math. Série 6 16, 427–475 (2007); URL

http://afst.cedram.org/item?id=AFST_2007_6_16_3_427_0.

[14] J.-P. Bourguignon, P. Gauduchon; Spineurs, opérateurs de Dirac et variations de métriques; Comm. Math. Phys. 144, 581–599 (1992).

[15] X. Dai, X. Wang, G. Wei; On the stability of Riemannian manifold with parallel spinors; Invent. Math. 161, 151–176 (2005); URL

http://dx.doi.org/10.1007/s00222-004-0424-x.

[16] J. M. Figueroa-O’Farrill; Breaking the M-Waves; Class. Quant. Grav. 17, 2925––2947 (2000).

[17] T. Friedrich; On the spinor representation of surfaces in Euclidean 3-space.; J. Geom. Phys. 28, 143–157 (1998).

[18] N. Ginoux; The Dirac spectrum; Lecture Notes in Mathematics, vol. 1976, Springer-Verlag, Berlin 2009; URL

http://dx.doi.org/10.1007/978-3-642-01570-0.

[19] N. Ginoux, G. Habib, S. Raulot; A new upper bound for the Dirac operators on hypersurfaces; Pacific J. Math. 278, 79–101 (2015); URL

http://dx.doi.org/10.2140/pjm.2015.278.79.

[20] R. Haslhofer, R. Müller; Dynamical stability and instability of Ricci-flat metrics; Math. Ann. 360, 547–553 (2014).

[21] A.  Hermann;  Dirac eigenspinors for generic met-

rics; Ph.D. thesis; Universität Regensburg (2012); URL

https://epub.uni-regensburg.de/25024/.

[22] A. Ikemakhen; Groupes d’holonomie et spineurs parallèles sur les variétés pseudo-riemanniennes complètement réductibles; C. R. Math. Acad. Sci. Paris 339, 203–208 (2004); URL

http://dx.doi.org/10.1016/j.crma.2004.04.029.

[23] K. Kröncke; Stability of Einstein metrics under Ricci flow; Comm. Anal. Geom. (to appear) 28 (2020).

[24] R. Kusner, N. Schmitt; The spinor representation of

Minimal surfaces;  Unpublished preprint (1995);  arXiv,

http://www.arxiv.org/abs/dg-ga/9512003.

[25] R. Kusner, N. Schmitt; The spinor representation of

surfaces in space;  Unpublished preprint (1996);  arXiv,

http://www.arxiv.org/abs/dg-ga/9610005.

[26] H. B. Lawson, M.-L. Michelsohn; Spin Geometry, Princeton University Press, Princeton 1989.

[27] T. Leistner; Lorentzian manifolds with parallel spinors; Ph.D. thesis; Humboldt-Universität Berlin (2003).

[28] T. Leistner; On the classification of Lorentzian holonomy groups; J. Differential Geom. 76, 423–484 (2007); URL

http://projecteuclid.org/euclid.jdg/1180135694.

[29] T. Leistner, A. Lischewski; Hyperbolic Evolution Equations, Lorentzian Holonomy, and Riemannian Generalised Killing Spinors; J. Geom. Anal. 29, 33–82 (2019); URL

http://dx.doi.org/10.1007/s12220-017-9941-x.

[30] A. Lischewski; The Cauchy problem for parallel spinors as first-order symmetric hyperbolic system (2015); preprint, arXiv:1503.04946.

[31] O. Müller, N. Nowaczyk; A universal spinor bundle and the Einstein-Dirac-Maxwell equation as a variational theory; Lett. Math. Phys. 107, 933–961

(2017); URL http://dx.doi.org/10.1007/s11005-016-0929-4.

[32] A. Murcia, C. Shabazi; Parallel spinors on globally hyperbolic Lorentzian four-manifolds (2020); preprint, arXiv:2011.02423.

[33] N. Nowaczyk; Continuity of Dirac spectra; Ann. Glob. Anal. Geom. 44,

541–563 (2013); URL http://dx.doi.org/10.1007/s10455-013-9381-1.

[34] H.-B. Rademacher; Generalized Killing spinors with imaginary Killing function and conformal Killing fields; in Global differential geometry and global analysis (Berlin, 1990); Lecture Notes in Math., vol. 1481; 192–198,

Springer, Berlin 1991; URL http://dx.doi.org/10.1007/BFb0083642.

[35] T. Schick, D. Wraith; Non-negative versus positive scalar curvature; J. Math. Pures Appl. (9) 146, 218–232 (2021); URL

http://dx.doi.org/10.1016/j.matpur.2020.09.010.

[36] J. Seipel; Cauchy problems on Lorentzian manifolds with parallel vector and spinor fields, Master thesis, Universität Regensburg, 2019.

[37] A. Trautman; Spinors and the Dirac operator on hypersurfaces. I. General theory; J. Math. Phys. 33, 4011–4019 (1992); URL

http://dx.doi.org/10.1063/1.529852.

[38] A. Trautman; The Dirac operator on hypersurfaces; Acta Phys. Polon. B 26, 1283–1310 (1995).

[39] M. Y. Wang; Preserving parallel spinors under metric deformations; Indiana Univ. Math. J. 40, 815–844 (1991); URL

http://dx.doi.org/10.1512/iumj.1991.40.40037.

[40] E. Witten; A new proof of the positive energy theorem; Commun. Math. Phys. 80, 381–402 (1981).